### v1

Abstract
We study belief polarization among Bayesian agents observing public information about a multidimensional state. [27] show that divergence in the sense of first-order stochastic dominance is impossible for one dimensional beliefs, but we find that in multidimensional settings it can occur for all marginal beliefs, even with infinitely many signals. At the same time, we extend their impossibility result: divergence in the sense of multidimensional stochastic dominance is impossible. For an intermediate stochastic order, polarization may arise only in the short-run. We provide necessary and sufficient conditions on signal structures for persistent polarization and discuss implications for polarization in actions.
1 Introduction
Polarization is a buzzword in current political discourse. Merriam-Webster named it its 2024 Word of the Year, noting that it is “one idea that both sides of the political spectrum agree on.” [^2] See Merriam-Webster “2024 Word of the Year: Polarization,” accessed June 26, 2025: merriam-webster.com/wordplay/word-of-the-year. [31] clarify that polarization can be defined both as a state—one in which society’s views are concentrated at opposite extremes, or as a process—i.e., the movement towards such a state. Furthermore, polarization can pertain to attitudes, beliefs, preferences etc. This paper focuses on beliefs and on polarization as a process. Specifically, we will use the following notion, based on the definition in [32] (henceforth DW).
Polarization occurs if the beliefs of two agents move further apart and in opposite directions after they observe the same information.
At first glance, this type of polarization appears at odds with rational Bayesian models. A common feature of such models is the prediction that when agents observe the same information their beliefs move closer together. Therefore, widespread polarization seemingly offers evidence of irrationality on at least one side. This intuition is formalized by several theoretical results, notably the impossibility theorem of [27] (henceforth BHK). Considering two Bayesian agents with different priors, they show that observing a common public signal cannot cause beliefs to diverge in the sense of first-order stochastic dominance. [^3] Meaning that if agent $H$’s prior stochastically dominates agent $L$’s prior, it cannot be the case that $H$’s posterior dominates $H$’s prior while $L$’s posterior is dominated by $L$’s prior. Interpreting polarization as irrational has deep implications. It fits the broader view that we are living in a post-truth era, marked by erosion of shared facts and a narrowing scope for constructive public discourse [41, 36].
In BHK, agents’ beliefs are defined over a one-dimensional state space. In practice, political beliefs are multidimensional, spanning foreign policy, the economy, cultural issues, and more. The first contribution of this paper is to show that once beliefs are multidimensional, a strong and practically relevant form of Bayesian polarization becomes possible: agents’ marginal beliefs can diverge simultaneously in the sense of stochastic dominance on each dimension. In other words, the type of divergence that [27] rule out in one dimension may arise coordinatewise across all dimensions.
The analysis also clarifies the limits of Bayesian polarization. Two distinctions matter: whether polarization is measured after a single signal realization (one-shot) or after an infinite sequence of i.i.d. realizations (limit); and which stochastic order is used to compare beliefs. The coordinatewise order described above requires stochastic dominance on each marginal dimension separately, and here both one-shot and limit polarization are possible. The paper’s second contribution is to establish the restrictions that Bayesian updating places on polarization, by showing impossibility results for stronger notions of divergence. The multidimensional stochastic dominance order is the standard generalization of one-dimensional stochastic dominance to higher dimensions. Under this order, both one-shot and limit polarization are impossible—an extension of BHK’s impossibility result to any number of dimensions. The upper-orthant order requires divergence on events involving multiple dimensions being simultaneously high (or low). It is stronger than the coordinatewise order but weaker than multidimensional stochastic dominance. [^4] Both the multidimensional stochastic order and the upper-orthant order have been suggested as generalizations of univariate stochastic dominance [48]. The former preserves the property that one distribution dominates another if and only if it dominates it in expectation for every increasing function. The latter is weaker, and preserves the property that one distribution dominates the other if and only if its distribution function dominates the other’s at every point. Under this intermediate criterion, one-shot polarization is possible but limit polarization is not. Table 1 summarizes.
Taken together, the results show that while there are forces pushing Bayesian agents towards agreement, these still allow for significant polarization in the face of common information. We demonstrate this idea from a different angle by considering polarization in beliefs on individual states. In any finite state space, there are always two states—the minimum- and maximum-likelihood states—on which both agents’ beliefs must move in the same direction. On all other states, however, their beliefs may move in opposite directions.
Mathematically, the BHK impossibility result generalizes, but substantively it is the coordinatewise order—not multidimensional stochastic dominance—that best reflects how polarization is measured in practice. Survey evidence often tracks divergence across separate issue dimensions such as climate change, trade, or cultural issues (see, e.g., [43]). The possibility result shows that this form of polarization is fully consistent with Bayesian updating, even in the long run. By contrast, the stronger orders capture divergence on combinations of issues, e.g., whether the state is simultaneously high on both economic and cultural dimensions. The impossibility results imply that agreement may persist on such joint events, even when marginals diverge. Evidence of polarization on such issue combinations would be harder to reconcile with rational updating.
Table 1: Possibility of Polarization
Stochastic Order One‐shot Polarization Limit Polarization
Coordinatewise $\checkmark$ $\checkmark$
Upper Orthant $\checkmark$ X
Stochastic Dominance X X
The third contribution is a complete characterization of the signals that can lead to limit coordinatewise polarization in two-dimensional state spaces. An immediate necessary condition is partial identification: there must be at least one additional state that is observationally equivalent to the true state of the world. Otherwise, beliefs converge to the truth, precluding polarization. [^5] Formally, if the signal is $X$, the state space is $\Theta$, and the true state is $\hat{\theta}$, then there must be at least one state $\theta\in\Theta\setminus\{\hat{\theta}\}$ such that for every realization $x$: $\Pr[X=x|\theta]=\Pr[X=x|\hat{\theta}$]. Otherwise, Doob’s Consistency Theorem ([33]) implies that, from any full-support prior, beliefs converge to a point mass on $\hat{\theta}$. Fixing the true state, any polarizing signal must therefore induce a set of states observationally equivalent to it, which we call the identified set. The problem of identifying polarizing signals thus reduces to characterizing the identified sets that permit polarization.
Under a mild technical condition, Theorem 5 shows that an identified set can lead to polarization if and only if it satisfies three geometric conditions. Both the set and its complement must be spanning—each must include states that attain the minimal and maximal values on every dimension. The set must be balanced—it cannot include all states that lie strictly below some point while excluding all of the states that lie weakly above it, or vice versa. And it must be non-compensatory—it cannot be the case that whenever states in the set are high on one dimension they are low on the other.
To illustrate the theorem, consider a simple $2\times 2$ state space $\Theta=\{x_{1},x_{2}\}\times\{y_{1},y_{2}\}$ with $x_{1}<x_{2}$ and $y_{1}<y_{2}$. Think of the $x$ dimension as describing the state of the economy, and the $y$ dimension the state of foreign policy. Suppose Democrats view both as low and Republicans view both as high. Two agents, $L$ and $H$, have full support priors over this state space. Agent $L$’s views are closer to the Democratic position: her prior is stochastically dominated by agent $H$’s on each dimension.
As shown in Figure 1, only two sets are spanning with spanning complements. These are the diagonal $\{(x_{1},y_{1}),(x_{2},y_{2})\}$ and the off-diagonal $\{(x_{1},y_{2}),(x_{2},y_{1})\}$. The off-diagonal violates the non-compensatory condition, since whenever a state is low on the economic dimension it is high on the foreign policy dimension, and vice-versa. The diagonal, on the other hand, satisfies all three conditions. Thus, a corollary of Theorem 5 is that in $2\times 2$ state spaces, limit coordinatewise polarization can arise if and only if the identified set is the diagonal. That is, polarization is possible only when the signal reveals that the two dimensions are perfectly correlated (and nothing else).
[…]
(a) set and its complement not spanning
[…]
(b) compensatory
[…]
(c) biased; complement not spanning
[…]
(d) uniquely satisfies the three conditions
Figure 1: Applying Theorem 5 to the $2\times 2$ space.
The result connects to a central stylized fact: the steady rise in polarization among US political elites since the 1970s has coincided with rising ideological consistency, with positions on one issue increasingly predicting positions on others [40]. The theorem’s conditions permit arbitrarily high correlation between the dimensions but the non-compensatory condition rules out strong negative correlation. [^6] In the $2\times 2$ case limit polarization can arise only when agents learn that the two dimensions are perfectly correlated. Generally, learning that the two dimensions are perfectly correlated can always lead to polarization while learning that they are perfectly negatively correlated cannot. This points to a complementary Bayesian mechanism to existing accounts of polarization: if elites encountered evidence of strong correlations among issues, their posteriors would have polarized while becoming more internally consistent, in line with the observed pattern. A suggestive illustration is the fall of ‘Third Way’ politics. This movement, championed in the 1990s by Bill Clinton and Tony Blair, combined conservative economic policies with progressive social stances. Its decline by the early 2000s [37] may have signaled the instability of such mixed positions, strengthening ideological consistency and thereby contributing to polarization.
The spanning and balance conditions imply that polarizing signals must (i) preserve the possibility that the state takes very low or very high values on each dimension, and (ii) avoid decisively indicating that the state is simultaneously high (or low) on both dimensions. This echoes journalistic norms of ‘balanced coverage’, which may seem counterintuitive, since media bias is commonly viewed as a key driver of polarization. Yet the biased-media account typically works through selective exposure, with each side consuming different evidence that supports its prior (see, e.g., [39]). By contrast, we study polarization in the presence of public information. In this setting, evidence that decisively supports one side cannot lead to polarization, but balanced evidence can. Historical accounts suggest that balanced coverage has, at times, prolonged disagreement. For instance, [45] document how tobacco industry leaders exploited the media’s adherence to the Fairness Doctrine to sustain controversy about the health risks of smoking, despite overwhelming scientific evidence.
While the analysis centers on beliefs, polarization is often discussed in terms of actions, such as voting, campaign donations, or social media posts. In the concluding section, we develop a simple political-economy model that connects the two. The key insight is that whether action polarization arises depends on the separability of preferences: robust action polarization arises when utility functions are additively separable across dimensions, but not when they display strong complementarities.
The paper proceeds as follows. Section 2 reviews related literature. Section 3 presents the model and the definitions of polarization, and Section 4 reviews preliminaries on stochastic orders and Bayesian learning. Section 5 develops the possibility and impossibility results, and Section 6 characterizes the identified sets that lead to limit coordinatewise polarization. Section 7 briefly discusses the probability of belief polarization, and polarization in actions.
2 Related Literature
Our notion of polarization builds on DW, who study a one-dimensional state space and define polarization as divergence in the stochastic dominance order after a single signal realization. DW show that polarization is impossible under Bayesian updating if signals satisfy the monotone likelihood ratio property (MLRP), and in particular in the standard linear-normal model. They also present examples suggesting that polarization may arise when MLRP fails, though, as later noted by BHK these examples display only divergence of expectations, not polarization in the stochastic dominance order. BHK adopt the DW definition and prove that polarization is impossible when agents are Bayesian and the state space is finite and one dimensional. [^7] The examples that DW present in detail do not display polarization according to their definition. They also sketch additional examples with continuum state spaces for which they do not specify the priors. Technically, these do not violate BHK, which assumes a discrete state space. Whether polarization is possible in continuum state spaces is an open question.
BHK’s impossibility result cemented the intuitive notion that belief polarization is inconsistent with Bayesian learning. This notion is in line with classic results in [33] and [29], which predict different forms of belief convergence among Bayesians learning from the same information. To reconcile the perceived theoretical impossibility with the apparent prevalence of polarization, several papers have introduced learning biases that generate polarization. These include ambiguity aversion (BHK), confirmation bias [46], misperceptions about selective sharing on a network [30], and correlation neglect [38].
Other work has challenged this view by showing that some forms of polarization are consistent with Bayesian updating. The closest to ours are [26] and [28], who demonstrate polarization in expectations on one dimension when agents hold beliefs over a two-dimensional state. [^8] Both focus on $2\times 2$ state spaces where the choice of stochastic order is irrelevant, but [26] note that their observation can be generalized to polarization in expectations.
[35] provide examples of Bayesian networks that may lead to polarization regarding the probability of a hypothesis. Put in our terms, this is another example of one-dimensional polarization in a two-dimensional state space. [47] show the possibility of belief dilation: for a family of prior beliefs, observing evidence may lead the upper and lower bounds for the probability of a given event to diverge. [25] study asymptotic disagreement rather than polarization. They focus on the magnitude of disagreement rather than its direction, showing that vanishing uncertainty about likelihood functions can lead to substantial asymptotic disagreement. [44] develop a model of persistent polarization due to rational inattention. [34] presents a philosophical account of predictable polarization, where rational agents’ beliefs diverge in directions they can anticipate ex-ante.
This paper generalizes earlier results on the possibility of Bayesian polarization by demonstrating the possibility of polarization in first order stochastic dominance on all dimensions simultaneously, for any number of dimensions, and following an infinite sequence of i.i.d. signals. It contributes to the debate on the rationality of belief polarization by showing both the possibility of Bayesian polarization and its limits. While most previous papers focused on short-term polarization, we distinguish between one-shot and asymptotic polarization and show that this distinction matters: under the upper orthant order, one-shot polarization is possible but asymptotic polarization is not. Finally, we characterize the signals that can lead to limit polarization and connect polarization in beliefs to polarization in actions.999[28] characterize the signals that lead to one-shot polarization in their setting, motivated by experimental evidence and focused on polarization in populations. Since their setting is more specific and the notions of polarization differ, the conditions are not directly comparable.
3 Model
Two agents, agent $L$ and agent $H$, are learning about the value of a parameter $\hat{\theta}\in\Theta\subset\mathbb{R}^{d}$, where $d\geq 2$. Each agent $i\in\{L,H\}$ has prior $P^{i}$ with full support on $\Theta$. This state space is a finite product space, $\Theta=\Theta_{1}\times\ldots\times\Theta_{d}$ where $1<|\Theta_{i}|<\infty$ for all $i\in[d]$. [^10] Throughout, for any integer $n\in\mathbb{N}$ we denote $n]=\{1,2,\ldots,n\}$.
Agents observe a realization $x\in\mathcal{X}$ of a public signal $X$, where $\mathcal{X}$ is finite. The likelihood function for this realization, known to both agents, is $\ell(\theta):=\Pr[X=x|\theta]$ for every state $\theta\in\Theta$. Upon observing the signal, the agents follow Bayes rule to update their beliefs to the respective posteriors $Q^{L}$ and $Q^{H}$. We assume that $\ell(\theta)>0$ for at least one state, so that Bayesian updating is well defined.
Definition 1 (One-shot polarization).
Let $\succeq$ be a partial order over $\Delta(\Theta)$. Say that one-shot $\succeq$-polarization occurs upon observing $x$ if
$Q^{L}\prec\ P^{L}\prec P^{H}\prec Q^{H}$
We will also consider the case where the agents observe an infinite sequence of i.i.d. signals $X_{1},X_{2},\ldots$, each with likelihood $\ell$ as above. Here, we let $Q^{L}(t),Q^{H}(t)$ denote the posterior after observing $X_{1},\ldots,X_{t}$, and let $Q^{L}(\infty),Q^{H}(\infty)$ denote the limit posteriors: $Q^{L}(\infty):=\lim_{t\rightarrow\infty}Q^{L}(t)$, and $Q^{H}:=\lim_{t\rightarrow\infty}Q^{H}(t)$ (we show below that there are unique limit distributions to which these sequences converge almost surely).
Definition 2 (Limit polarization).
Let $\succeq$ be a partial order over $\Delta(\Theta)$. Say that limit $\succeq$-polarization occurs if the limit posteriors satisfy
$Q^{L}(\infty)\prec\ P^{L}\prec P^{H}\prec Q^{H}(\infty)$
4 Preliminaries
4.1 Stochastic Orders
For two vectors $x,y\in\mathbb{R}^{d}$, say $x\leq y$ if $x_{i}\leq y_{i}$ for all $i\in[d]$, and say $x\ll y$ if $x_{i}<y_{i}$ for all $i\in[d]$. For a set $\Gamma\in\mathbb{R}^{d}$ define $\min(\Gamma)=\{\,x\in\Gamma\mid\nexists y\in\Gamma\setminus\{x\}\;\text{s.t.}\;y\leq x\,\}$. Similarly, $\max(\Gamma)$ is the subset of vectors in $\Gamma$ such that no other vector in $\Gamma$ is weakly above them. We will be considering finite sets for which the min and max sets are always non empty. Since $\Theta$ is a finite product space its min and max sets are singletons. Let $\underaccent{\bar}{\theta}$ denote the unique minimal element in $\Theta$ and $\bar{\theta}$ denote the unique maximal element. For general subsets $\Gamma\subset\Theta$, $\min(\Gamma)$ and $\max(\Gamma)$ may contain more than one element.
We now define three classes of sets, which we will use to define three stochastic orders. A set $U\subset\Theta$ is an upper set if $\theta\in U$ whenever $\theta\in\Theta$, $\theta\geq\delta$ and $\delta\in U$. A set $U\subset\Theta$ is an upper orthant if it is of the form $\{\,\theta\in\Theta\mid\theta_{1}>a_{1},\theta_{2}>a_{2},\ldots,\theta_{d}>a_{d}\,\}$ for some $a\in\mathbb{R}^{d}$. A set $U\subset\Theta$ is an upper projection if it is of the form $\{\,\theta\in\Theta\mid\theta_{i}>a\,\}$ for some $a\in\mathbb{R}$ and $i\in[d]$. [^11] Upper sets and upper orthants are standard definitions in the literature on stochastic orders, upper projection is not [48]. Note that we could have equivalently defined each of these classes of sets over $\mathbb{R}^{d}$, in which case we would have said that a set $U\subset\Theta$ is an upper set if there exists an upper set $\tilde{U}\in\mathbb{R}^{d}$ such that $U=\Theta\cap\tilde{U}$.
Definition 3 (Stochastic orders).
Let $P^{L},P^{H}$ be two probability measures over $\Theta$ and $P^{L}_{i},P^{H}_{i}$ be their respective marginals on the $i$ coordinate. Say:
  1. 1.
    $P^{L}$ is stochastically dominated by $P^{H}$, denoted $P^{L}\preceq_{st}P^{H}$ if $P^{L}(U)\leq P^{H}(U)$ for all upper sets $U\subset\Theta$.
  2. 2.
    $P^{L}$ is upper orthant dominated by $P^{H}$, denoted $P^{L}\preceq_{uo}P^{H}$ if $P^{L}(U)\leq P^{H}(U)$ for all upper orthants $U\subset\Theta$.
  3. 3.
    $P^{L}$ is coordinatewise dominated by $P^{H}$, denoted $P^{L}\preceq_{cw}P^{H}$ if $P^{L}(U)\leq P^{H}(U)$ for all upper projections $U\subset\Theta$. Equivalently, $P^{L}_{i}\preceq_{st}P^{H}_{i}$ for all $i\in[d]$.
It is immediate that every upper projection is an upper orthant, but not vice versa, and that every upper orthant is an upper set, but not vice versa. This proves the following result, which states that the orders in Definition 3 are ordered from the strongest, or most restrictive, to the weakest. [^12] This relationship is also noted by [48] who do not formally define coordinatewise polarization but mention that upper orthant polarization implies univariate polarization on each coordinate.
Lemma 1.
$P^{L}\preceq_{st}P^{H}\Longrightarrow P^{L}\preceq_{uo}P^{H}\Longrightarrow P^{L}\preceq_{cw}P^{H}$ and none of the reverse implications hold.
Another way to define stochastic orders is via inequalities of expectations over a class of functions. Given a set of functions $\mathcal{U}\subset\mathbb{R}^{\Theta}$, and a partial order $\succeq$ over $\Delta(\Theta)$, say that $\succeq$ is generated by $\mathcal{U}$ if
$P^{L}\preceq P^{H}\iff E^{L}[u]\leq E^{H}[u]\quad\forall u\in\mathcal{U}$
Here $E^{L}[u]$ and $E^{H}[u]$ are the expectations of $u$ according to the distributions $P^{L}$ and $P^{H}$, respectively. We will use this notation throughout the paper. Table 2 below summarizes the generating functions for the three stochastic orders defined in 3. Here, by increasing function we mean a function $u:\Theta\rightarrow\mathbb{R}$ that is increasing according to the order $\geq$. A product of nonnegative univariate increasing functions is a function $u:\Theta\rightarrow\mathbb{R}$ such that $u(\theta)=\Pi_{i=1}^{d}u_{i}(\theta_{i})$ where each $u_{i}$ is nonnegative and increasing, and a sum of nonnegative univariate increasing functions is defined similarly. The generating functions for the orders $\succeq_{st}$ and $\succeq_{uo}$ are presented in [48], and the generating functions for $\succeq_{cw}$ are derived in Lemma 2 below. [^13] By linearity of the expectation, if an order $\succeq$ is generated by a set $\mathcal{U}$ of functions then it is also generated by the set of all functions that are convex combinations of functions in $\mathcal{U}$.
Table 2: Generating Functions
Stochastic Order Generating Functions
$\preceq_{st}$ Increasing functions
$\preceq_{uo}$ Products of nonnegative
univariate increasing functions
$\preceq_{cw}$ Sums of univariate increasing functions
Lemma 2.
The order $\preceq_{cw}$ is generated by the set
$\mathcal{U}=\{\,u\in\mathbb{R}^{\Theta}\mid u(\theta)=\sum_{i=1}^{d}u_{i}(\theta_{i})\text{ where }u_{i}:\Theta_{i}\rightarrow\mathbb{R}\text{ are increasing}\,\}$
4.2 Limit Posteriors
Consider an agent with prior $P\in\Delta(\Theta)$ observing an infinite sequence $X_{1},X_{2},\ldots$ of i.i.d. signals. Let $\ell$ be the likelihood function of each of these signals, with $\ell(x;\theta)=\Pr[X=x|\theta]$. [^14] Outside this section we fix the signal realization $x$ and simply write $\ell(\theta)$. Let $\hat{\theta}$ be the true state of the world, the identified set is
$\hat{\Gamma}:=\{\,\theta\in\Theta\mid\ell(x;\theta)=\ell(x;\hat{\theta})\,\forall x\in\mathcal{X}\,\}$
Let $Q(t)$ be the posterior upon observing all signals up to and including $X_{t}$ and define $Q^{*}\in\Delta(\Theta)$ as:
$Q^{*}(\theta)=\begin{cases}\frac{P(\theta)}{P(\hat{\Gamma})},&\theta\in\hat{\Gamma}\\ 0,&\text{else}\end{cases}$
Lemma 3.
For every $\theta\in\Theta$, $Q(t)(\theta)\rightarrow Q^{*}(\theta)$ almost surely.
Lemma 3 implies that the limit belief following an infinite sequence of signals is almost surely equal to the posterior belief that would arise upon observing a single signal that revealed that the state is in the identified set and nothing else. Formally, say that $X$ is a partitional signal if there exists a partition $\mathcal{G}$ of $\Theta$ and a one-to-one mapping $x:\mathcal{G}\rightarrow\mathcal{X}$ such that for every $\theta\in G\in\mathcal{G}$:
$\ell(x(G);\theta)=1\text{ and }\ell(x^{\prime};\theta)=0\text{ for all }x^{\prime}\neq x(G)$
The signal $X$ reveals what partition cell the state lies in, and two states in the same partition cell are observationally equivalent.
Henceforth, we identify partitional signal realizations with their associated partition cells, i.e., we set $\mathcal{X}=\mathcal{G}$ and $x(\Gamma)=\Gamma$ for all $\Gamma\in\mathcal{G}$. Lemma 3 implies that limit polarization is equivalent to one-shot polarization upon observing a partitional signal. That is, it shows that the following definition is equivalent to Definition 2.
Definition 2’.
Fix a partitional signal $X$ with partition $\mathcal{G}$. Let $P^{L},P^{H}\in\Delta(\Theta)$ be the agents’ priors and let $Q^{L},Q^{H}$ be their respective posteriors upon observing a realization $\Gamma$ of $X$. Say that limit polarization occurs for the identified set $\Gamma$ if
$Q^{L}\prec\ P^{L}\prec P^{H}\prec Q^{H}$
This formulation will be our working definition. One advantage is that it clarifies that limit polarization is a special case of one-shot polarization. An immediate implication is that if limit polarization is possible then so is one-shot polarization. Conversely, if one shot-polarization is impossible then limit polarization is impossible as well.
5 (Im)Possibility of Polarization
This section presents the results about the possibility and impossibility of polarization that are summarized in Table 1. We begin with the coordinatewise order.
Theorem 1.
Limit $\succeq_{cw}$-polarization is possible.
We show in Section 6 that many identified sets can lead to coordinatewise polarization. The simplest, which works in any dimension, is the set $\Gamma=\{\underaccent{\bar}{\theta},\bar{\theta}\}$ that includes just the minimal and maximal states. If agent $L$’s prior significantly favors $\underaccent{\bar}{\theta}$ over $\bar{\theta}$ and is uniform over the remaining states, and agent $H$’s is the mirror image, then this set may lead to coordinatewise polarization. The following example illustrates in a $2\times 2$ state space and the proof of Theorem 1 generalizes to any finite product space $\Theta$.
Example 1.
Let $\Theta=\{x_{1},x_{2}\}\times\{y_{1},y_{2}\}$ with the priors given in Figure 2.
[…]
Figure 2: Priors for Example 1.
Polarization occurs for the identified set $\Gamma=\{(x_{1},y_{1}),(x_{2},y_{2})\}$, because the marginals on any coordinate $i$ satisfy:
$Q^{L}_{i}=(3/4,1/4)\prec_{st}P^{L}_{i}=(5/8,3/8)\prec_{st}P^{H}_{i}=(3/8,5/8)\prec_{st}Q^{H}_{i}=(1/4,3/4)$
We saw above that if limit polarization is possible according to a stochastic order then one-shot polarization according to that order is possible as well. Hence, an immediate corollary of Theorem 1 is that one-shot coordinatewise polarization is possible, thereby completing the first row of Table 1.
We now turn to the upper orthant order. First, consider a $2\times 2$ state space as in Example 1. This state space has four upper orthants, illustrated in Figure 3. Note that the priors and identified set $\Gamma$ presented in Example 1 do not lead to upper orthant polarization, since for both agents the mass on the singleton upper set $U_{4}$ increases after observing the signal. Hence $Q^{L}$ is not upper orthant dominated by her $P^{L}$ so upper orthant polarization does not occur, and this will the be the case for any identified set that includes $\bar{\theta}=(x_{2},y_{2})$. On the other hand, if the identified set does not include $\bar{\theta}$ then for both agents the mass on $U_{4}$ decreases after observing the signal. But this implies that $P^{H}$ is not upper orthant dominated by $Q^{H}$ so that upper orthant polarization does not occur. This argument extends to any state space $\Theta$ and leads to the following result.
[…]
(a) $U_{1}$
[…]
(b) $U_{2}$
[…]
(c) $U_{3}$
[…]
(d) $U_{4}$
Figure 3: The four upper orthants in a $2\times 2$ state space.
Theorem 2.
Limit $\succeq_{uo}$-polarization is impossible.
With partitional signals, agents’ posteriors place more mass than their priors on any state in the identified set, and zero mass outside. Thus, for any singleton event, agents’ beliefs move in the same direction. This was not an issue for coordinatewise polarization—where none of the relevant events (upper projections) are singletons—but it rules out limit upper orthant polarization.
Without the restriction that the signal is partitional, agents’ beliefs may move in opposite directions also over singleton events, and this allows for one-shot upper orthant polarization.
Theorem 3.
One-shot $\succeq_{uo}$-polarization is possible.
The proof builds on the observation that the posterior probability of a state exceeds its prior if and only if its likelihood is larger than the expected likelihood under the prior. Formally, if $P$ is the prior and $Q$ the posterior after observing a realization with likelihood function $\ell$, then for any $\theta$, $Q(\theta)>P(\theta)$ iff $\ell(\theta)>E_{P}[\ell]$. Lemma 9 in the appendix gives a more general version.
To see how this enables upper orthant polarization, consider the $2\times 2$ space, and a signal realization $x$ with the likelihood function $\ell$ presented in Figure 4. If agent $L$’s prior concentrates around the minimal state $\underaccent{\bar}{\theta}$ then her expected likelihood over the state space will be close to $\ell(\underaccent{\bar}{\theta})=1$. Sine this exceeds $\ell(\bar{\theta})=\frac{1}{2}$, her belief on $\bar{\theta}$ decreases after observing $x$. Conversely, we can construct agent $H$’s prior so that her expected likelihood is below $\frac{1}{2}$, leading her belief on $\bar{\theta}$ to increase. Thus, beliefs diverge on the singleton upper orthant. With this likelihood function and suitable priors, beliefs may diverge on all upper orthants.
Note that unlike the symmetric likelihood function that generated coordinatewise polarization, here a key feature is that the likelihoods of $\underaccent{\bar}{\theta}$ and $\bar{\theta}$ differ. This allows for agents to initially update in opposite directions but means that the state is identified so agents’ limit beliefs converge to the true state.
[…]
Figure 4: Likelihood function for upper orthant polarization.
We complete this section by showing that one-shot polarization according to the stochastic dominance order is impossible. An immediate corollary is that limit polarization according to this order is also impossible, thereby completing the last row of Table 2.
Theorem 4.
One-shot $\succeq_{st}$-polarization is impossible.
Theorem 4 extends Theorem 1 of BHK from finite one-dimensional state spaces to any finite multidimensional state space. [^15] The proof does not rely on the assumption that $\Theta$ is a product space. The proof follows theirs but instead of using induction on the states as in BHK, the induction argument is on the sizes of upper sets from $1$ to $|\Theta|$.
To get an intuition for why one-shot upper orthant polarization is possible while one-shot stochastic polarization is not, recall that all upper orthants are upper sets but not vice-versa. In the $2\times 2$ state space there is one upper set that is not an upper orthant, the “upper L” presented in Figure 5. For polarization to occur the high agent’s beliefs must increase on the upper set $\{\bar{\theta}\}$ and decrease on $\{\underaccent{\bar}{\theta}\}$, which is the complement of the upper $L$. The opposite must occur for the low agent, but this is violates general restriction implied by Bayes rule.
[…]
Figure 5: Upper set that is not an upper orthant.
The second claim in Lemma 4 presents this restriction: If agent $L$ updates down on a state where agent $H$ updates up, then it cannot also be the case that $L$ updates up on a state on a state where $H$ updates down. This observation is also implicitly used by BHK. The first claim is of independent interest, as it conveys the key message of this section in the context of general (unordered) state spaces: Bayesian updating imposes only mild restrictions on belief polarization, allowing beliefs to diverge on all but two states. [^16] Note that updating in opposite directions is not synonymous with polarization as agents may update in opposite directions towards each other. But it is straightforward to construct an example where agents’ beliefs move away from each other on all but two states.
Lemma 4.
Let $P,P^{\prime}\in\Delta(\Theta)$ be full support priors over a finite state space $\Theta$ and $Q,Q^{\prime}$ be the corresponding posteriors following signal realization $x$. If beliefs move at all then:
  1. 1.
    There exist two states—the minimum and maximum likelihood states—on which beliefs move in the same direction. On all other states, beliefs may move in opposite directions.
  2. 2.
    If beliefs move in opposite directions, the divergence is consistent across states: If there exists a state $\theta\in\Theta$ for which $Q(\theta)-P(\theta)\leq 0\leq Q^{\prime}(\theta)-P^{\prime}(\theta)$, with one of these inequalities strict, then there does not exist a state $\theta^{\prime}$ at which $Q(\theta^{\prime})-P(\theta^{\prime})\geq 0\geq Q^{\prime}(\theta^{\prime})-P^{\prime}(\theta^{\prime})$.
Proof. Recall that $\ell(\theta):=\Pr[X=x|\theta]$ is the likelihood function. Let $\theta^{-}\in\arg\min(\ell)$ and $\theta^{+}\in\arg\max(\ell)$ be a min- and a max-likelihood state. Let $E_{P}[\ell],E_{P^{\prime}}[\ell]$ denote the expectations of $\ell$ according to $P$ and $P^{\prime}$, respectively. By Lemma 9 in the appendix, for every $\theta\in\Theta$ we have $Q(\theta)>P(\theta)$ if and only if $\ell(\theta)>E_{P}[\ell]$, and similarly for $Q^{\prime},P^{\prime}$. For beliefs to move at all, the likelihood function must not be constant, which implies $\ell(\theta^{-})<\ell(\theta^{+})$. Thus, since $P$ and $P^{\prime}$ have full support, $\ell(\theta^{-})<\min\{E_{P}[\ell],E_{P^{\prime}}[\ell]\}$ and $\ell(\theta^{+})>\max\{E_{P}[\ell],E_{P^{\prime}}[\ell]\}$. So beliefs move in the same direction on $\theta^{-},\theta^{+}$.
To see that beliefs on all other states may move in opposite directions, assume that $\ell(\theta^{-})=a$, $\ell(\theta^{+})=c$ and $\ell(\theta)=b$ for all other $\theta\in\Theta$, where $a<b<c$. It is easy to construct full-support priors with $E_{P}[\ell]>b>E_{P^{\prime}}[\ell]$. For such priors, for every state $\theta\in\Theta\setminus\{\theta^{-},\theta^{+}\}$ we have $Q(\theta)-P(\theta)<0<Q^{\prime}(\theta)-P^{\prime}(\theta)$.
Finally, assume that there exists $\theta\in\Theta$ for which $Q(\theta)-P(\theta)<0\leq Q^{\prime}(\theta)-P^{\prime}(\theta)$ (here the first inequality is the strict one w.l.o.g). Then, $E_{P}[\ell]>\ell(\theta)\geq E_{P^{\prime}}[\ell]$. So, there does not exist $\theta^{\prime}\in\Theta$ at which $Q(\theta^{\prime})-P(\theta^{\prime})\geq 0\geq Q^{\prime}(\theta^{\prime})-P^{\prime}(\theta^{\prime})$, as that would imply $E_{P}[\ell]\leq E_{P^{\prime}}[\ell]$, a contradiction.
Lemma 4 highlights two restrictions on belief divergence in the face of public information. Yet agents can still polarize through two channels. The first is state-level polarization: beliefs can diverge on all but two states. The second is event-level polarization: even when beliefs move in the same directions on every state, they may diverge on events because the magnitude of belief updates depends on the priors. For one agent, small decreases on some states within an event may be outweighed by large increases on others. For the other, the opposite may hold, leading to polarization on the event.
In the long run, the first channel disappears: for any two full-support priors, limit posteriors move in the same direction on every state, placing positive mass on the identified set and zero elsewhere. Hence, limit coordinatewise polarization operates through the second channel. By contrast, upper orthant and stochastic polarization require divergence on at least one individual state. This is impossible for limit beliefs, but in the short run the combination of the first and second channels allows for upper orthant polarization.
6 Conditions for Coordinatewise Polarization
This section characterizes the information that can lead to limit coordinatewise polarization in two-dimensional state spaces. Since limit beliefs are pinned down by the priors and the identified set, the problem reduces to specifying conditions on the identified set. We now define the relevant conditions.
Definition 4 (Spanning).
A set $\Gamma\subset\Theta$ is spanning (in $\Theta$) if it obtains the lowest and highest values on each coordinate. Formally, for every coordinate $i$ there exist $\theta,\theta^{\prime}\in\Gamma$ such that $\theta_{i}=\underaccent{\bar}{\theta}_{i},\theta^{\prime}_{i}=\bar{\theta}_{i}$.
Definition 5 (Biased and balanced).
$\Gamma\subset\Theta$ is biased downward if there exists $\theta\in\Theta$, with $\theta\gg\underaccent{\bar}{\theta}$, such that $\Gamma$ contains all states strictly below $\theta$ and none of the states weakly above it. [^17] Formally, $\{\theta^{\prime}\in\Theta|\theta^{\prime}\ll\theta\}\subset\Gamma$ and $\{\theta^{\prime}\in\Theta|\theta^{\prime}\geq\theta\}\cap\Gamma=\emptyset.$ $\Gamma\subset\Theta$ is biased upward if there exists $\theta\in\Theta$, with $\theta\ll\bar{\theta}$ such that $\Gamma$ contains all states strictly above $\theta$ and none of the states weakly below it. $\Gamma$ is balanced if it is not biased downward or upward.
Definition 6 (Compensatory).
$\Gamma\subset\Theta$ is compensatory if there exists $\theta\in\Theta$, with $\theta\ll\bar{\theta}$ such that every state in $\Gamma$ is neither weakly below $\theta$ nor strictly above it. That is, for every $\theta^{\prime}\in\Gamma$ neither $\theta^{\prime}\leq\theta$ nor $\theta^{\prime}\gg\theta$ holds.181818 Equivalently, there exists $\hat{\theta}\in\Theta$ with $\hat{\theta}\gg\underaccent{\bar}{\theta}$ such that for every $\theta^{\prime}\in\Gamma$ neither $\theta^{\prime}\geq\hat{\theta}$ nor $\theta^{\prime}\ll\hat{\theta}$ holds. If the condition as it appears in the text is satisfied for $\theta\ll\underaccent{\bar}{\theta}$ then the version in this footnote is satisfied for $\hat{\theta}=\min\{\theta^{\prime}\in\Theta|\theta^{\prime}\gg\theta\}$.
In the characterization theorem below, the relevant condition is that a set be non-compensatory.
The theorem also requires one technical condition. We first introduce a strong form of coordinatewise dominance that rules out equalities between the marginal cdfs. Let $P^{L}$ and $P^{H}$ be two distributions whose supports are contained in $X=\{x_{1},\ldots,x_{n}\}\subset\mathbb{R}$ where $x_{1}<x_{2}<\ldots<x_{n}$. Let $F^{L}$ and $F^{H}$ denote their cdfs. Say that $P^{L}$ is strongly stochastically dominated by $P^{H}$ on $X$, denoted $P^{L}\prec^{*}_{st}P^{H}$ if $F^{L}(x_{i})>F^{H}(x_{i})$ for all $1\leq i<n$. For two distributions $P^{L},P^{H}$ whose supports are contained in $\Theta\subset\mathbb{R}^{d}$ say that $P^{L}$ is strongly coordinatewise dominated by $P^{H}$ on $\Theta$, denoted $P^{L}\prec_{cw}^{*}P^{H}$ if $P^{L}_{i}\prec^{*}_{st}P^{H}_{i}$ on $\Theta_{i}$ for all $i\in[d]$.
We now define a strong form of limit coordinatewise polarization, which requires strong dominance between the priors. Let $P^{L},P^{H}\in\Delta(\Theta)$ be priors and let $Q^{L},Q^{H}\in\Delta(\Theta)$ be the posteriors upon observing a partitional signal with realization $\Gamma$. Say that strong coordinatewise polarization occurs if
$Q^{L}\prec_{cw}P^{L}\prec^{*}_{cw}P^{H}\prec_{cw}Q^{H}$
Say that an identified set $\Gamma$ can strongly coordinatewise polarize if there exist priors for which strong coordinatewise polarization occurs for a partitional signal with realization $\Gamma$.
We are now ready to state the characterization theorem. Here and below we let $\Gamma^{C}=\Theta\setminus\Gamma$ denote the complement of $\Gamma$ in $\Theta$.
Theorem 5.
Let $\Theta\subset\mathbb{R}^{2}$. An identified set $\Gamma\subset\Theta$ can strongly coordinatewise polarize if and only if all of the following hold: (i) $\Gamma$ and $\Gamma^{C}$ are spanning, (ii) $\Gamma$ is balanced, (iii) $\Gamma$ is non-compensatory.
The conditions of Theorem 5 can be expressed geometrically. A set is spanning if it intersects every extreme row and column. A set is biased downward if it is contained in a “lower L,” while its complement is contained in the corresponding “upper L.” That is, there exists a state such that, for the Euclidean coordinate system originating at that state, $\Gamma$ covers the lower left quadrant, while $\Gamma^{C}$ covers the upper right quadrant. Similarly, a set is biased upward if it covers the upper right quadrant while its complement covers the lower left quadrant. Finally, a set is compensatory if it is contained in an off-diagonal, i.e., in the union of the upper left and lower right quadrants originating at some state.
As illustrated in Figure 1, in a $2\times 2$ state space, the unique set that satisfies all three conditions is the diagonal $\Gamma=\{(x_{1},y_{1}),(x_{2},y_{2})\}$. In general two-dimensional state spaces, any set that contains the extreme states $\underaccent{\bar}{\theta}$ and $\bar{\theta}$ is spanning and satisfies conditions (ii) and (iii), so can polarize as long its complement is spanning, i.e., as long as it does not cover any extreme row or column. In particular, in symmetric state spaces, the diagonal can always polarize while the off-diagonal cannot polarize. However, there are sets that satisfy the conditions and do not contain one or both of the extreme states. For example, in any two-dimensional product space greater than $2\times 2$ the set $\Theta\setminus\{\underaccent{\bar}{\theta},\bar{\theta}\}$ that includes all states except the two extremes satisfies all three conditions.
We now turn to the proof of Theorem 5, which requires the following result.
Lemma 5.
Let $P$ be a full support prior over $\Theta$ and $Q$ the posterior after receiving a set signal with realization $\Gamma$. Let $P|_{\Gamma},{P}|_{\Gamma^{C}}$ be the conditional distributions of $P$ over $\Gamma$ and $\Gamma^{C}$ respectively. Then $Q\preceq_{cw}P\iff P|_{\Gamma}\preceq_{cw}{P}|_{\Gamma^{C}}$.
If an identified set violates one of the first two conditions of Theorem 5 then either the low agent’s beliefs cannot move down in the coordinatewise order or the high agent’s beliefs cannot move up. For example, if $\Gamma$ is biased downward, then it cannot be the case that $P^{H}\prec_{cw}Q^{H}$. If an identified set satisfies the first two conditions but violates the third then, while it is possible that the low agent’s beliefs move down, and possible that the high agent’s beliefs move up, it cannot be the case that both occur for the same identified set.
To illustrate, consider again the $2\times 2$ state space. Here the only non-compensatory set that satisfies the first two conditions is the “off-diagonal” $\Gamma=\{(x_{1},y_{2}),(x_{2},y_{1})\}$. Assume towards contradiction that $\Gamma$ can polarize. By Lemma 6 and the fact that the coordinatewise order is generated by sums of univariate increasing functions, for any such function $u$ it must be the case that $E^{L}[u|\Gamma]\leq E^{L}[u|\Gamma^{C}]$ and $E^{H}[u|\Gamma]\geq E^{H}[u|\Gamma^{C}]$. Consider the function $u(\theta)=u_{1}(\theta_{1})+u_{2}(\theta_{2})$, where $u_{1}(x_{1})=u_{2}(y_{1})=0$, and $u_{1}(x_{2})=u_{2}(y_{2})=1$. The values of $u$ over all states in $\Theta$ are summarized in Figure 6. Since $u(\theta)=1$ for all $\theta\in\Gamma$, we get $E^{L}[u|\Gamma]=E^{H}[u|\Gamma]=1$. So the above implies that $E^{L}[u|\Gamma^{C}]\geq 1$ and $E^{H}[u|\Gamma^{C}]\leq 1$. But then the unconditional expectations satisfy $E^{L}[u]\geq 1\geq E^{H}[u]$. However, the assumption $P^{L}\prec^{*}_{cw}P^{H}$ implies $E^{L}[u]<E^{H}[u]$, a contradiction. [^19] The final inequality follows from Lemma 11 in the appendix. The reason we need to assume that $P^{H}$ strongly dominates $P^{L}$ is to ensure that this inequality is strict.
[…]
Figure 6: Values of $u(\theta)$. States in the off-diagonal $\Gamma$ are outlined in red.
The proof that the non-compensatory condition is necessary is a straightforward extension of this example and is relegated to the appendix along with the rest of the necessity direction. The sufficiency proof is more involved. Its outline is presented here, with intermediate lemmas proved in the appendix. The key step is Lemma 6, which is proved by induction on the size of the state space. Although the sufficiency proof does not provide the polarizing priors explicitly, the induction argument of Lemma 6 can be adapted to construct them recursively.
Proof of Theorem 5 (sufficiency). Since $\Theta\subset\mathbb{R}^{2}$ we can assume $\Theta=\{x_{1},\ldots,x_{m}\}\times\{y_{1},\ldots,y_{n}\}$ for $m,n\geq 2$. Throughout this proof, let $F^{L}_{i},F^{H}_{i}$ denote the cdfs of the marginal distributions of $P^{L}$ and $P^{H}$, respectively on the $i$ coordinate. Let $F^{L}_{i}|_{\Gamma},F^{L}_{i}|_{\Gamma}^{C}$ denote the cdfs of the marginal distributions of $P^{L}|_{\Gamma}$ and $P^{L}|_{\Gamma^{C}}$, and similarly define $F^{H}_{i}|_{\Gamma},F^{H}_{i}|_{\Gamma^{C}}$.
The conditions are sufficient:
Let $\Gamma\subset\Theta$ be a set that satisfies all three conditions. We will establish the existence of $P^{L},P^{H}\in\Delta(\Theta)$ such that strong coordinatewise polarization occurs upon observing $\Gamma$. The proof proceeds in three steps.
Step 1: A method for constructing strongly coordinatewise ordered distributions.
Recall that a subset of a partially ordered set is an antichain if any two distinct elements in the set are incomparable to each other. We will consider antichains that are subsets of $\Theta$, under the order $\leq$ defined above. We first define a partial order over such antichains.
Definition 7.
Given two nonempty antichains $A,B\subset\Theta\subset\mathbb{R}^{2}$ say that $B$ antichain dominates $A$ in $\Theta$, denoted $B>^{\Theta}_{AC}A$ if the following hold:
  1. 1.
    AA and $B$ are disjoint.
  2. 2.
    AA hits the minimal value on both coordinates (i.e., there exist $\theta,\theta^{\prime}\in A$ such that $\theta_{1}=\underaccent{\bar}{\theta}_{1}$, $\theta^{\prime}_{2}=\underaccent{\bar}{\theta}_{2}$) and $B$ hits the maximal value on both coordinates.
  3. 3.
    $A\cup B$ is non-compensatory.
Lemma 6.
Let $A,B\subset\Theta$ be two nonempty antichains such that $B>^{\Theta}_{AC}A$. Then, there exist distributions $P_{A}$ with full support on $A$ and $P_{B}$ with full support on $B$ such that $P_{A}\prec^{*}_{cw}P_{B}$.
Step 2: Applying the method to construct intermediate distributions.
We now apply Lemma 6 to construct distributions $P^{L}|_{\min(\Gamma)},P^{L}|_{\max(\Gamma^{C})},P^{H}|_{\max(\Gamma)}$ and $P^{H}|_{\min(\Gamma^{C})}$, each supported on the set indicated in its subscript, such that the following conditions hold: [^20] We have not yet defined the distributions $P^{L}$ and $P^{H}$ but will define them in the next step so that they are consistent with this notation. For example, $P^{L}|_{\min(\Gamma)}$ will be the conditional distribution of $P^{L}$ on the set $\min(\Gamma)$.
$\begin{cases}P^{L}|_{\min(\Gamma)}\prec^{*}_{cw}P^{H}|_{\max(\Gamma)}\\ P^{L}|_{\min(\Gamma)}\prec^{*}_{cw}P^{L}|_{\max(\Gamma^{C})}\\ P^{H}|_{\min(\Gamma^{C})}\prec^{*}_{cw}P^{H}|_{\max(\Gamma)}\end{cases}$ (1)
By Lemma 6, to verify that such distributions exist, it suffices to prove the following.
Lemma 7.
Under the conditions of Theorem 5 all of the following hold:
  1. 1.
    $\min(\Gamma)<^{\Theta}_{AC}\max(\Gamma)$
  2. 2.
    $\min(\Gamma)<^{\Theta}_{AC}\max(\Gamma^{C})$
  3. 3.
    $\min(\Gamma^{C})<^{\Theta}_{AC}\max(\Gamma)$
Step 3: Constructing the priors $P^{L}$ and $P^{H}$.
Let $P^{L}|_{\min(\Gamma)},P^{L}|_{\max(\Gamma^{C})},P^{H}|_{\max(\Gamma)},P^{H}|_{\min(\Gamma^{C})}$ be distributions that satisfy the conditions in (1). Recall that $\Theta=\{x_{1},\ldots,x_{m}\}\times\{y_{1},\ldots,y_{n}\}$. By strong coordinatewise dominance, there exists $\varepsilon>0$ such that for all $i\in[m-1]$ and $j\in[n-1]$,
$\begin{split}F^{L}_{1}|_{\min(\Gamma)}(x_{i})&>\max\{F^{H}_{1}|_{\max(\Gamma)}(x_{i}),F^{L}_{1}|_{\max(\Gamma^{C})}(x_{i})\}+\varepsilon\\ F^{L}_{2}|_{\min(\Gamma)}(y_{j})&>\max\{F^{H}_{2}|_{\max(\Gamma)}(y_{j}),F^{L}_{2}|_{\max(\Gamma^{C})}(y_{j})\}+\varepsilon\end{split}$
For any set $A\subset\Theta$, let $U_{A}$ denote the uniform distribution over $A$, and let $F^{U}_{i}$ denote its marginal cdf on the $i$ coordinate. Let $\delta\in(0,1)$ and define [^21] These distributions are defined as weighted mixtures of their summands. For example, for every $\theta\in\Theta$: $P^{L}(\theta)=(1-\delta)\left[(1-\delta)P^{L}|_{\min(\Gamma)}(\theta)+\delta U_{\Gamma\setminus\min(\Gamma)}(\theta)\right]+\delta\left[(1-\delta)P^{L}|_{\max(\Gamma^{C})}(\theta)+\delta U_{\Gamma^{C}\setminus\max(\Gamma^{C})}(\theta)\right].$
[…]
With these definitions of the priors, the following lemma establishes polarization, completing the sufficiency proof.
Lemma 8.
Polarization occurs for the identified set $\Gamma$ and the priors in (2).
7 Final Remarks
This paper focused on the possibility of belief polarization among Bayesian agents and on conditions for it to arise. We conclude by briefly discussing two additional issues: the probability of polarization, and the relationship between polarization of beliefs and polarization of actions.
7.1 Probability of Polarization
The following example demonstrates that limit coordinatewise polarization is not only possible but can also be very probable. However, there is a tradeoff between the probability of polarization and its magnitude.
Example 2.
Let $\delta>0$. Consider the $2\times 2$ state space $\Theta=\{x_{1},x_{2}\}\times\{y_{1},y_{2}\}$ with the priors given in Figure 7.
[…]
Figure 7: Priors for Example 2
The posteriors upon observing a partitional signal with identified set $\Gamma=\{(x_{1},y_{1}),(x_{2},y_{2})\}$ are given in Figure 8.
[…]
Figure 8: Posteriors for Example 2
In this symmetric example, a natural measure for the magnitude of polarization as a function of $\delta$ is the difference between the post- and pre-signal gaps in the agents’ marginal beliefs:
$M(\delta)=Q^{L}_{1}(x_{1})-Q^{H}_{1}(x_{1})-(P^{L}_{1}(x_{1})-P^{H}_{1}(x_{1}))=Q^{L}_{2}(y_{1})-Q^{H}_{2}(y_{1})-(P^{L}_{2}(y_{1})-P^{H}_{2}(y_{1}))$
A simple computation shows $M(\delta)=\frac{\delta}{2}$ so the magnitude of polarization is linearly increasing in $\delta$.
Note that both agents agree on the prior probability that the state is in $\Gamma$ and that polarization only occurs if the true state is in $\Gamma$. [^22] This is true regardless of how we define the partitional signal for states outside of $\Gamma$. The signal can either reveal the true state or only reveal that the state is not in $\Gamma$ and both cases will not lead to polarization. This probability is $P^{L}(\Gamma)=P^{H}(\Gamma)=1-\delta$, which is linearly decreasing in $\delta$. So, in this example there is a direct tradeoff between the probability of polarization and its magnitude. When $\delta$ is close to 0, polarization almost surely occurs (according to both agents’ subjective beliefs) but the magnitude when it occurs is very small. As $\delta$ increases, polarization becomes less likely, but its magnitude increases.
7.2 Action Polarization
So far we have focused on belief polarization, but in practice polarization is most often observed in actions. We now demonstrate how the results for belief polarization translate into results for action polarization, using a simple political-economy model.
Consider a finite population of citizens in the run-up to an election between two candidates, an incumbent (I) and a challenger (C). Each citizen $i$ has a utility function $u_{i}:\{I,C\}\times\Theta\rightarrow\mathbb{R}$ where $u_{i}(n,\theta)$ denotes citizen $i$’s future payoff in the event that candidate $n$ is elected in state $\theta$. Assume there is no uncertainty about the incumbent’s future performance, so $\theta\in\Theta\subset\mathbb{R}^{d}$ only describes the challenger’s future performance across $d$ different policy dimensions. That is, for each citizen, $u_{i}(I,\theta)$ is constant in $\theta$, while $u_{i}(C,\theta)$ is not.
The population is partitioned into two groups $L$ and $H$, with beliefs $P^{L},P^{H}\in\Delta(\Theta)$. Think of these groups as representing partisans for two opposing parties. Each citizen evaluates their expected utility under the challenger, $U_{i}(C)$, according to their belief and utility function, and takes an action $a_{i}\in\mathcal{A}\subset\mathbb{R}$ accordingly. For instance, if $i\in L$ then $U_{i}(C)=E^{L}[u_{i}(C,\theta)]$. Assume that actions are ordered from against the challenger to in favor of, and that agents with higher $U_{i}(C)$ take higher actions. For example, a low action in $\mathcal{A}$ might correspond to donating to the incumbent’s campaign, a low-moderate action to tweeting in their favor, and a high action to actively campaigning for the challenger.
Assume group $H$ is initially taking higher actions, i.e., has a more favorable view of the challenger. Say action polarization occurs if a piece of public information moves the actions of all $L$ citizens down and the actions of all $H$ citizens up.
To see the relationship between belief and action polarization, consider first the coordinatewise order, which is generated by sums of univariate increasing functions. If agents $i\in L$ and $j\in H$ both have utility functions that are additively separable and increasing in each dimension, then a signal that induces coordinatewise belief polarization will also drive their expectations apart, resulting in action polarization. By Theorem 1, coordinatewise belief polarization is possible both in the short-run and in the long-run limit. Hence, if all agents’ utility functions are additively separable and increasing, then action polarization is possible as well.
The implications of impossibility results on belief polarization are a bit more subtle. Although one-shot belief polarization is impossible according to the stochastic order, which is generated by increasing functions, this does not imply that one-shot action polarization is impossible when agents’ utility functions are increasing. Rather, it implies that for any priors and signal, there exists at least one increasing utility function $u$ such that if an $L$ agent and an $H$ agent share this utility function then their actions will not polarize. Thus, impossibility of belief polarization rules out the following stronger notion of action polarization.
For a set $\mathcal{U}$ of utility functions $u:\Theta\rightarrow\mathbb{R}$, say that $\mathcal{U}$-action polarization occurs if there exist priors and a signal realization such that action polarization occurs whenever all agents’ utility functions are in $\mathcal{U}$. There is an immediate mapping from the results on belief polarization summarized in Table 1 to results on $\mathcal{U}$-action polarization according to the generating classes of functions. The latter are summarized in Table 3.
Table 3: Possibility of $\mathcal{U}$-action polarization
$\mathcal{U}$ One-shot Polarization Limit Polarization
Sums of univariate
increasing functions
$\checkmark$ $\checkmark$
Products of nonnegative
univariate
increasing functions
$\checkmark$ X
Increasing functions X X
Table 3 shows that separable utility functions allow for a broader scope of action polarization than utility functions that display strong complementarities across dimensions. Intuitively, the impossibility of one-shot belief polarization according to the stochastic order implies that, following any informative signal, there is always some upper set on which beliefs converge rather than polarize. A utility function that displays strong complementarities may assign arbitrarily high values within this set and arbitrarily low values outside it, thereby ruling out divergence in expectations.
By Theorem 5, for coordinatewise polarization to arise, the public signal must convey some positive association between the different dimensions. This informational condition is counterbalanced by a restriction on preferences: strong complementarities across dimensions may preclude polarization in actions. Thus, action polarization requires information that ties different dimensions together and preferences that are sufficiently separable across them.
Appendix A Proofs
Proof of Lemma 2. Let $P^{L},P^{H}\in\Delta(\Theta)$ such that $P^{L}\preceq_{cw}P^{H}$. Let $u=\sum_{i}u_{i}\in\mathcal{U}$. Since $P^{L}_{i}\preceq_{st}P^{H}_{i}$ for all $i\in[d]$, we also have $E_{P^{L}_{i}}[u_{i}]\leq E_{P^{H}_{i}}[u_{i}]$ for all $i\in[d]$. So,
$E^{L}[u]=\sum_{i=1}^{d}E_{P^{L}_{i}}[u_{i}]\leq\sum_{i=1}^{d}E_{P^{H}_{i}}[u_{i}]=E^{H}[u]$
For the other direction, assume that $E^{L}[u]\leq E^{H}[u]$ for all $u\in\mathcal{U}$. Let $j\in[d]$ and let $u_{j}:\Theta_{j}\rightarrow\mathbb{R}$ be an increasing function. Define $u_{i}\equiv 0$ for all $i\neq j$ and $u:=\sum_{i=1}^{d}u_{i}$. Then $u(\theta)=u_{j}(\theta_{j})$ for all $\theta$ so $E^{L}[u]\leq E^{H}[u]$ implies $E_{P^{L}_{j}}[u_{j}]\leq E_{P^{H}_{j}}[u_{j}]$. Since this holds for all $j\in[d]$ and for any increasing $u_{j}:\Theta_{j}\rightarrow\mathbb{R}$ we get $P^{L}_{j}\preceq_{st}P^{H}_{j}$ for all $j\in[d]$, so $P^{L}\preceq_{cw}P^{H}$.
Proof of Lemma 3. Let $\mathcal{G}=\{\Gamma_{1},\Gamma_{2},\ldots,\Gamma_{n}\}$ be a partition of $\Theta$ according to the equivalence relation $\ell(\cdot;\theta)\equiv\ell(\cdot;\theta^{\prime})$. Note that $\hat{\theta}\in\hat{\Gamma}\in\mathcal{G}$. For any $\Gamma\in\mathcal{G}$ let $\ell(\cdot;\Gamma)$ be the common likelihood function for states in $\Gamma$. Then, for any $\Gamma\in\mathcal{G}$, if the realizations of the first $t$ signals are $x_{1},\ldots,x_{t}$,
$Q(t)(\Gamma)=\sum_{\theta\in\Gamma}\frac{P(\theta)\Pi_{\tau=1}^{t}\ell(x_{\tau};\theta)}{\sum_{\theta^{\prime}\in\Theta}P(\theta^{\prime})\Pi_{\tau=1}^{t}\ell(x_{\tau};\theta^{\prime})}=\frac{P(\Gamma)\Pi_{\tau=1}^{t}\ell(x_{\tau};\Gamma)}{\sum_{\Gamma^{\prime}\in\mathcal{G}}P(\Gamma^{\prime})\Pi_{\tau=1}^{t}\ell(x_{\tau};\Gamma^{\prime})}$
That is, the agent’s posteriors over $\mathcal{G}$ are exactly as they would be in a learning problem where the state space is $\mathcal{G}$, the true state is $\hat{\Gamma}$, and the prior and likelihood function for each state $\Gamma$ are $P(\Gamma)$ and $\ell(\cdot;\Gamma)$, respectively. Note that for this state space it is the case that $\Gamma\neq\Gamma^{\prime}\Longrightarrow\ell(\cdot;\Gamma)\neq\ell(\cdot;\Gamma^{\prime})$ (i.e., there exists $x\in\mathcal{X}$ such that $\ell(x;\Gamma)\neq\ell(x;\Gamma^{\prime})$). Thus, by Doob’s Theorem for posterior consistency, $\lim_{t\rightarrow\infty}Q(t)(\hat{\Gamma})=1\text{ a.s}$. [^23] The original Theorem is [33], here we apply [42] Theorem 2.4. The theorem has some technical assumptions that are easily satisfied by our finite state space. The key assumption is that the state is identifiable: $\Gamma\neq\Gamma^{\prime}\Longrightarrow\ell(\cdot;\Gamma)\neq\ell(\cdot;\Gamma^{\prime})$. This implies that for any $\theta\notin\Gamma$, $\lim_{t\rightarrow\infty}Q(t)(\theta)=0=Q^{*}(\theta)$ a.s. Finally, for any $\theta\in\Gamma$, following any sequence $x_{1},\ldots,x_{t}$ of signal realizations,
$\frac{Q(t)(\theta)}{Q(t)(\Gamma)}=\frac{P(\theta)\Pi_{\tau=1}^{t}\ell(x_{\tau};\Gamma)}{P(\Gamma)\Pi_{\tau=1}^{t}\ell(x_{\tau};\Gamma)}=\frac{P(\theta)}{P(\Gamma)}$
So, for any $t$, $Q(t)(\theta)=Q(t)(\Gamma)\frac{P(\theta)}{P(\Gamma)}$, which implies that $\lim_{t\rightarrow\infty}Q(t)(\theta)=\frac{P(\theta)}{P(\Gamma)}$ a.s.
Proof of Theorem 1. Let $\varepsilon\in(0,1)$, set $\Gamma=\{\underaccent{\bar}{\theta},\bar{\theta}\}$, and define
$P^{L}(\underaccent{\bar}{\theta})=\frac{1}{|\Theta|}+\frac{\varepsilon}{|\Theta|},\,P^{L}(\bar{\theta})=\frac{1}{|\Theta|}-\frac{\varepsilon}{|\Theta|},\text{ and }P^{L}(\theta)=\frac{1}{|\Theta|}\,\,\forall\theta\notin\Gamma$
Define $P^{H}$ as a mirror image of $P^{L}$,
$P^{H}(\underaccent{\bar}{\theta})=\frac{1}{|\Theta|}-\frac{\varepsilon}{|\Theta|},\,P^{H}(\bar{\theta})=\frac{1}{|\Theta|}+\frac{\varepsilon}{|\Theta|},\text{ and }P^{H}(\theta)=\frac{1}{|\Theta|}\,\,\forall\theta\notin\Gamma$
Let $Q^{L}$ and $Q^{H}$ be the agents’ posteriors after learning that the identified set is $\Gamma$:
$\begin{split}Q^{L}(\underaccent{\bar}{\theta})&=\frac{1+\varepsilon}{2},\,Q^{L}(\bar{\theta})=\frac{1-\varepsilon}{2},\text{ and }Q^{L}(\theta)=0\,\,\forall\theta\notin\Gamma\\ Q^{H}(\underaccent{\bar}{\theta})&=\frac{1-\varepsilon}{2},\,Q^{H}(\bar{\theta})=\frac{1+\varepsilon}{2},\text{ and }Q^{H}(\theta)=0\,\,\forall\theta\notin\Gamma\end{split}$
Let $i\in[d]$ be a coordinate, and let $n:=|\Theta_{i}|$ denote the number of elements on this coordinate. Then, the prior and posterior marginals on the $i$ coordinate are:
$\begin{split}\quad P^{L}_{i}&=\left(\frac{1}{n}+\frac{\varepsilon}{|\Theta|},\frac{1}{n},\ldots,\frac{1}{n},\frac{1}{n}-\frac{\varepsilon}{|\Theta|}\right),\quad Q^{L}_{i}=\left(\frac{1+\varepsilon}{2},0,\ldots,0,\frac{1-\varepsilon}{2}\right)\\ P^{H}_{i}&=\left(\frac{1}{n}-\frac{\varepsilon}{|\Theta|},\frac{1}{n},\ldots,\frac{1}{n},\frac{1}{n}+\frac{\varepsilon}{|\Theta|}\right),\quad Q^{H}_{i}=\left(\frac{1-\varepsilon}{2},0,\ldots,0,\frac{1+\varepsilon}{2}\right)\end{split}$
Note that $P^{L}_{i}\prec_{st}P^{H}_{i}$ for any $\varepsilon>0$, and that both $Q^{L}_{i}\prec_{st}P^{L}_{i}$ and $P^{H}_{i}\prec_{st}Q^{H}_{i}$ hold if and only if $\varepsilon\geq\frac{|\Theta|(n-2)}{n(|\Theta|-2)}$ (recall that we assume $n\geq 2$ and $|\Theta|>2$). The right-hand side of this inequality is strictly less than $1$ since $|\Theta|\geq 2n>n$. Hence, there exists $\varepsilon\in(0,1)$ such that coordinatewise polarization occurs.
Proof of Theorem 2. Assume towards contradiction that limit $\succeq_{uo}$-polarization occurs for an identified set $\Gamma\subset\Theta$. Then $\Gamma\neq\Theta$, so for any $\theta\in\Gamma$ we have $Q^{L}(\theta)>P^{L}(\theta)$ and $Q^{H}(\theta)>P^{H}(\theta)$. The set $\{\bar{\theta}\}$ is an upper orthant in $\Theta$. If $\bar{\theta}\in\Gamma$ then $Q^{L}(\bar{\theta})>P^{L}(\bar{\theta})$, a contradiction to $Q^{L}\prec_{uo}P^{L}$. Conversely, if $\bar{\theta}\notin\Gamma$ then $Q^{H}(\bar{\theta})=0<P^{H}(\bar{\theta})$, a contradiction to $P^{H}\prec_{uo}Q^{H}$.
Proof of Theorem 3. Let $n\in\mathbb{N}$ and $\varepsilon\in(0,1)$. Define priors as follows:
$\begin{split}P^{L}(\underaccent{\bar}{\theta})&=1-\frac{1}{n}-\frac{1}{n^{2}},P^{L}(\bar{\theta})=\frac{1}{n^{2}},\text{ and }P^{L}(\theta)=\frac{1}{n\left(|\Theta|-2\right)}\quad\forall\theta\notin\{\underaccent{\bar}{\theta},\bar{\theta}\}\\ P^{H}(\underaccent{\bar}{\theta})&=\frac{1}{n^{2}},P^{H}(\bar{\theta})=1-\frac{1}{n}-\frac{1}{n^{2}},\text{ and }P^{H}(\theta)=\frac{1}{n\left(|\Theta|-2\right)}\quad\forall\theta\notin\{\underaccent{\bar}{\theta},\bar{\theta}\}\end{split}$
Let $x$ be signal realization with likelihood function given by,
$\ell(\underaccent{\bar}{\theta})=1,\ell(\bar{\theta})=1-\varepsilon,\text{ and }\ell(\theta)=0\quad\forall\theta\notin\{\underaccent{\bar}{\theta},\bar{\theta}\}$
Let $Q^{L}$ and $Q^{H}$ be the posteriors after observing the realization $x$. We will show upper orthant polarization occurs for sufficiently large $n$. We first show that $P^{L}\prec_{uo}P^{H}$. Let $U\subsetneq\Theta$ be an upper orthant. Then $\bar{\theta}\in U$ and $\underaccent{\bar}{\theta}\notin U$. So, if $|U|=k$ then, for $n>2$,
$P^{L}(U)=\frac{1}{n^{2}}+\frac{k}{n\left(|\Theta|-2\right)}<1-\frac{1}{n}-\frac{1}{n^{2}}+\frac{k}{n\left(|\Theta|-2\right)}=P^{H}(U)$
So, $P^{L}\prec_{uo}P^{H}$.
We now show that $P^{L}\prec_{uo}Q^{L}$ by proving that for sufficiently large $n$, $P^{L}(U)<Q^{L}(U)$ for any upper orthant $U\subsetneq\Theta$. By Lemma 9 below, this is equivalent to proving that for sufficiently large $n$ $E^{L}[\ell|U]<E^{L}[\ell]$ for any upper orthant $U\subsetneq\Theta$. Let $U\subsetneq\Theta$ be an upper orthant and let $k=|U|$. Then $\bar{\theta}\in U$ and $\underaccent{\bar}{\theta}\notin U$ imply,
$E^{L}[\ell|U]=\frac{1-\varepsilon}{1+\frac{n(k-1)}{|\Theta|-2}}\leq 1-\varepsilon$
On the other hand, as $n\rightarrow\infty$
$E^{L}[\ell]=\left(1-\frac{1}{n}-\frac{1}{n^{2}}\right)\cdot 1+\frac{1}{n^{2}}\cdot(1-\varepsilon)+\frac{1}{n}\cdot 0\rightarrow 1$
Finally, we prove that for sufficiently large $n$, $P^{H}(U)<Q^{H}(U)$ for any upper orthant $U\subsetneq\Theta$. Again, let $U\subsetneq\Theta$ be an upper orthant and let $k=|U|$. As $n\rightarrow\infty$,
$E^{H}[\ell|U]=\left(1-\frac{1}{n}-\frac{1}{n^{2}}+\frac{k-1}{n\left(|\Theta|-2\right)}\right)^{-1}\left(\left(1-\frac{1}{n}-\frac{1}{n^{2}}\right)\cdot(1-\varepsilon)\right)\rightarrow 1-\varepsilon$
Since $E^{H}[\ell]$ is a weighted average of $E^{H}[\ell|U]$ and $E^{H}[\ell|\Theta\setminus U]$ to prove that $E^{H}[\ell|U]>E^{H}[\ell]$ it suffices to prove that $E_{P^{H}}[\ell|U]>E_{P^{H}}[\ell|\Theta\setminus U]$. And this holds for sufficiently large $n$, since as $n\rightarrow\infty$,
$E^{H}[\ell|\Theta\setminus U]=\frac{1}{1+\frac{n\left(|\Theta|-k-1\right)}{\left(|\Theta|-2\right)}}\rightarrow 0$
For the limit, we used the fact that $|U|=k<|\Theta|-1$. To see this, first note that for any upper orthant $U\subsetneq\Theta$ it must be the case that $\underaccent{\bar}{\theta}\notin U$. So, if $|U|\geq|\Theta|-1$, then $U=\Theta\setminus\{\underaccent{\bar}{\theta}\}$. Let $a\in\mathbb{R}^{d}$ be the origin of the upper orthant $U$, i.e. the point for which $U=\{\theta\in\Theta:\theta_{1}>a_{1},\theta_{2}>a_{2},\ldots,\theta_{d}>a_{d}\}$. Let $i\in[d]$. Since $U=\Theta\setminus\{\underaccent{\bar}{\theta}\}$ there exists a state $\theta\in U$ with $\theta_{i}=\underaccent{\bar}{\theta}_{i}$. Thus $a_{i}\leq\theta_{i}$. Since this holds for all $i\in[d]$ we have $a\leq\underaccent{\bar}{\theta}$, which implies $\underaccent{\bar}{\theta}\in U$, a contradiction.
Lemma 9.
Let $P\in\Delta(\Theta)$ be a prior belief with full support over $\Theta$ and $Q\in\Delta(\Theta)$ be the corresponding posterior after observing a realization $x$ of signal $X$. Then, for every $A\subset\Theta$: $Q(A)>P(A)$ if and only if $E_{P}[\ell|A]>E_{P}[\ell]$.
Proof of Lemma 9. Let $A\subset\Theta$. Since $Q$ is obtained by Bayesian updating with prior $P$ and likelihood $\ell$ we have,
$\begin{split}&Q(A)>P(A)\iff\frac{\sum_{\theta\in A}P(\theta)\ell(\theta)}{\sum_{\theta\in\Theta}P(\theta)\ell(\theta)}>P(A)\iff\frac{\sum_{\theta\in A}P(\theta)\ell(\theta)}{P(A)}>\sum_{\theta\in\Theta}P(\theta)\ell(\theta)\\ &\iff E_{P}[\ell|A]>E_{P}[\ell]\end{split}$
Here we used the fact that $P$ has full support and that $\ell(\theta)>0$ for at least one $\theta\in\Theta$ to multiply both sides of the inequality by $\frac{\sum_{\theta\in\Theta}P(\theta)\ell(\theta)}{P(A)}$.
Proof of Theorem 4. Assume towards contradiction that there exists a signal realization $x$ that leads to one shot $\succeq_{st}$-polarization. We will prove by induction on $n$ that for every upper set $U\subset\Theta$ of size $n$: $P^{L}(U)=Q^{L}(U)$ and $P^{H}(U)=Q^{H}(U)$. This contradicts the definition of $\succeq_{st}$-polarization, which requires the existence of upper sets $U$ and $U^{\prime}$ such that $P^{L}(U)>Q^{L}(U)$ and $P^{H}(U^{\prime})<Q^{H}(U^{\prime})$.
Base Case: Since $\{\bar{\theta}\}$ and $\Theta\setminus\{\underaccent{\bar}{\theta}\}$ are upper sets, the assumption $Q^{L}\prec_{st}P^{L}$ implies that $Q^{L}(\bar{\theta})\leq P^{L}(\bar{\theta})$ and $Q^{L}(\underaccent{\bar}{\theta})\geq P^{L}(\underaccent{\bar}{\theta})$. Similarly, the assumption $Q^{H}\succ_{st}P^{H}$ implies $Q^{H}(\bar{\theta})\geq P^{H}(\bar{\theta})$ and $Q^{H}(\underaccent{\bar}{\theta})\leq P^{H}(\underaccent{\bar}{\theta})$. By Lemma 4, none of these inequalities can be strict, so $Q^{L}(\bar{\theta})=P^{L}(\bar{\theta})$ and $P^{H}(\bar{\theta})=Q^{H}(\bar{\theta})$. Since $\{\bar{\theta}\}$ is the unique upper set of size $1$ this completes the base case. [^24] This theorem also holds without the assumption that $\Theta$ is a product space, in which case it may not have a unique maximum. The only change is in the base case, where we would let $\bar{\theta}$ be some maximal element of $\Theta$ and $\underaccent{\bar}{\theta}$ be some minimal element of $\Theta$.
Inductive Step: Let $1<n\leq|\Theta|$, and assume that the inductive hypothesis holds for all $1\leq n^{\prime}<n$. Consider an upper set $U$ of size $n$. Let $\theta_{U}$ be a minimal element of $U$ and define $A=U\setminus\{\theta_{U}\}$. Then $A$ is an upper set of size n1n-1. The assumption that $Q^{L}\prec_{st}P^{L}$ implies $P^{L}(A)+P^{L}(\theta_{U})=P^{L}(U)\geq Q^{L}(U)=Q^{L}(A)+Q^{L}(\theta_{U})$. The inductive hypothesis implies $P^{L}(A)=Q^{L}(A)$, which implies $P^{L}(\theta_{U})\geq Q^{L}(\theta_{U})$. By an analogous argument $P^{H}(\theta_{U})\leq Q^{H}(\theta_{U})$. So, as in the base case, Lemma 4 implies $P^{L}(\theta_{U})=Q^{L}(\theta_{U})$ and $P^{H}(\theta_{U})=Q^{H}(\theta_{U})$. Hence, $P^{L}(U)=Q^{L}(U)$ and $P^{H}(U)=Q^{H}(U)$, which completes the induction argument.
Proof of Lemma 5. For any $\theta\in\Gamma$: $Q(\theta)-P(\theta)=\frac{P(\theta)}{P(\Gamma)}-P(\theta)=\frac{P(\theta)}{P(\Gamma)}P(\Gamma^{C})$. So, for any coordinate $i\in[d]$ and $x\in\mathbb{R}$,
$\begin{split}&Q(\theta_{i}\leq x)-P(\theta_{i}\leq x)=\sum_{\theta\in\Gamma;\theta_{i}\leq x}\left[P(\theta)+\frac{P(\theta)}{P(\Gamma)}P(\Gamma^{C})\right]-\sum_{\theta\in\Theta;\theta_{i}\leq x}P(\theta)=\\ &\frac{P(\Gamma^{C})}{P(\Gamma)}\sum_{\theta\in\Gamma;\theta_{i}\leq x}P(\theta)-\sum_{\theta\in\Gamma^{C};\theta_{i}\leq x}P(\theta)=\frac{P(\Gamma^{C})}{P(\Gamma)}P(\Gamma\cap\{\theta_{i}\leq x\})-P(\Gamma^{C}\cap\{\theta_{i}\leq x\})\end{split}$
So,
$\begin{split}&Q\preceq_{cw}P\iff Q(\theta_{i}\leq x)\geq P(\theta_{i}\leq x)\quad\forall i\in[d],x\in\mathbb{R}\iff\\ &\frac{P(\Gamma\cap\{\theta_{i}\leq x\})}{P(\Gamma)}\geq\frac{P(\Gamma^{C}\cap\{\theta_{i}\leq x\})}{P(\Gamma^{C})}\quad\forall i\in[d],x\in\mathbb{R}\iff P|_{\Gamma}\preceq_{cw}P|_{\Gamma^{C}}\end{split}$
Rest of Proof of Theorem 5.
The conditions are necessary:
  1. (i)
    $\Gamma$ cannot polarize if it is not spanning or if $\Gamma^{C}$ is not spanning: Assume that $\Gamma$ is not spanning and assume towards contradiction that coordinatewise polarization occurs. If $\Gamma$ does not obtain the minimal value on one of the coordinates, w.l.o.g the first, then $F^{L}_{1}|_{\Gamma}(x_{1})=0<F^{L}_{1}|_{\Gamma^{C}}(x_{1})$. By Lemma 5, this violates $Q^{L}\prec_{cw}P^{L}$, a contradiction. Similarly, if $\Gamma$ does not obtain the maximal value on the first coordinate then $F^{H}_{1}|_{\Gamma}(x_{n-1})=1<F^{H}_{1}|_{\Gamma^{C}}(x_{n-1})$, a contradiction to $P^{H}\prec_{cw}Q^{H}$. Similar arguments show that $\Gamma^{C}$ must be spanning as well.
  2. (ii)
    $\Gamma$ cannot polarize if it is not balanced: Assume w.l.o.g. that $\Gamma$ is biased upward and assume towards contradiction that coordinatewise polarization occurs. Let $\theta^{*}\in\Theta$ be a state such that $\theta^{*}\ll\bar{\theta}$ and such that $\Gamma$ includes all states strictly above $\theta^{*}$ and none of the states weakly below it. Define the function $u:\Theta\rightarrow\mathbb{R}$ as $u(\theta)=u_{1}(\theta_{1})+u_{2}(\theta_{2})$ where
    $u_{1}(x)=\begin{cases}0&x\leq\theta^{*}_{1}\\ 1&x>\theta^{*}_{1}\end{cases};\quad u_{2}(y)=\begin{cases}0&y\leq\theta^{*}_{2}\\ 1&y>\theta^{*}_{2}\end{cases}$
    Note that $u(\theta)=0$ for all $\theta\leq\theta^{*}$, $u(\theta)=2$ for all $\theta\gg\theta^{*}$, and $u(\theta)=1$ otherwise. So, the assumption that $\Gamma$ does not contain any state weakly below $\theta^{*}$ and contains all states strictly above it implies that $E^{L}[u|\Gamma]>1$ and $E^{L}[u|\Gamma^{C}]<1$. [^25] Recall that these are the conditional expectations according to the prior $P^{L}$. Since $\theta^{*}\ll\bar{\theta}$ there exists states $\theta\in\Theta$ with $\theta\gg\theta^{*}$. So, the assumption that $P^{L}$ has full support implies the strict inequalities. Since $u$ is a sum of univariate increasing functions, and the coordinatewise order is generated by such functions, this violates $P^{L}|_{\Gamma}\prec_{cw}P^{L}|_{\Gamma^{C}}$, and therefore by Lemma 5 violates $Q^{L}\prec_{cw}P^{L}$, a contradiction.
  3. (iii)
    $\Gamma$ cannot polarize if it is compensatory: Assume that $\Gamma$ is compensatory and assume towards contradiction that coordinatewise polarization occurs. Let $\theta^{*}\in\Theta$ be such that $\theta^{*}\ll\bar{\theta}$ and such that for every $\theta^{\prime}\in\Gamma$ it is not the case that $\theta^{\prime}\leq\theta^{*}$ and it is not the case that $\theta^{\prime}\gg\theta^{*}$. Then, for the function $u$ defined in the previous step we have $u(\theta)\equiv 1$ for every $\theta\in\Gamma$. Next, by Lemma 11 below, the strong coordinatewise order implies strict inequality of expectations for nontrivial sums of univariate increasing functions. That is, $P^{L}\prec^{*}_{cw}P^{H}$ and the fact that $u$ is not constant over $\Theta$ imply that $E^{L}[u]<E^{H}[u]$. By Lemma 5, $Q^{L}\prec P^{L}$ implies $P^{L}|_{\Gamma}\prec_{cw}P^{L}|_{\Gamma^{C}}$, which implies
    $1=E^{L}[u|\Gamma]\leq E^{L}[u|\Gamma^{C}]$
    Therefore,
    $E^{L}[u]=P^{L}(\Gamma)E^{L}[u|\Gamma]+P^{L}(\Gamma^{C})E^{L}[u|\Gamma^{C}]\geq 1$
    Similarly, $P^{H}\prec_{cw}Q^{H}$ implies
    $1=E^{H}[u|\Gamma]\geq E^{H}[u|\Gamma^{C}]$
    which implies $E^{H}[u]\leq 1$. So we have $E^{L}[u]\geq 1\geq E^{H}[u]$, a contradiction.
Lemma 10.
Let $P^{L}$ and $P^{H}$ be two univariate distributions with a common finite support $X=\{x_{1},\ldots,x_{n}\}\subset\mathbb{R}$ where $x_{1}<x_{2}<\ldots<x_{n}$. Then $P^{L}\prec^{*}_{st}P^{H}$ implies that $E^{L}[u]<E^{H}[u]$ for any function $u:X\rightarrow\mathbb{R}$ that is weakly increasing and not constant.
Proof. Let $F^{L}$ and $F^{H}$ be the corresponding cdfs. Using summation by parts,
$\begin{split}E^{H}[u]-E^{L}[u]&=\sum_{i=1}^{n}u(x_{i})\left(P^{H}(x_{i})-P^{L}(x_{i})\right)\\ &=u(x_{n})\left(F^{H}(x_{n})-F^{L}(x_{n})\right)-\sum_{i=1}^{n-1}\left(F^{H}(x_{i})-F^{L}(x_{i})\right)\left(u(x_{i+1})-u(x_{i})\right)\\ &=\sum_{i=1}^{n-1}\left(F^{L}(x_{i})-F^{H}(x_{i})\right)\left(u(x_{i+1})-u(x_{i})\right)\end{split}$
Since $P^{L}\prec^{*}_{st}P^{H}$, we have $F^{L}(x_{i})-F^{H}(x_{i})>0$ for every $i\in[n-1]$. This, together with the fact that $u$ is weakly increasing and not constant implies that the summands in the final expression are weakly greater than 0 for all $i\in[n-1]$ and strictly greater than 0 for some $i\in[n-1]$, which implies $E^{H}[u]-E^{L}[u]>0$.
Lemma 11.
Let $P^{L},P^{H}\in\Delta(\Theta)$ be such that $P^{L}\prec^{*}_{cw}P^{H}$. Let $u:\Theta\rightarrow\mathbb{R}$, $u=\sum_{i}u_{i}$, be a sum of univariate increasing functions and assume that $u$ is not constant over $\Theta$. Then $E^{L}[u]<E^{H}[u]$.
Proof. Since $u$ is not constant, one of the functions $u_{i}$ must not be constant. Thus, since $P^{L}_{i}\prec^{*}_{st}P^{H}_{i}$ for all $i\in[d]$, Lemma 10 implies that $E_{P^{L}_{i}}[u_{i}]\leq E_{P^{H}_{i}}[u_{i}]$ for all $i\in[d]$ and $E_{P^{L}_{j}}[u_{j}]<E_{P^{H}_{j}}[u_{j}]$ for at least one $j\in[d]$ . So,
$E^{L}[u]=\sum_{i=1}^{d}E_{P^{L}_{i}}[u_{i}]<\sum_{i=1}^{d}E_{P^{H}_{i}}[u_{i}]=E^{H}[u]$
Proof of Lemma 6. For the proof it will be easier to denote $P^{A}$ and $P^{B}$ rather than $P_{A},P_{B}$ so that $P^{A}_{i}$ denotes the marginal of $P^{A}$ on the $i$ coordinate, $F^{A}_{i}$ denotes the cdf of $P^{A}_{i}$, and likewise for $B$.
We prove by induction on $s=|\Theta|$. Note that for $\Theta$ to be a two-dimensional product set its cardinality $s$ must be a composite number so the base case is $s=4$ and the induction is over all composite numbers $s$.
Base case: Assume $\Theta=\{x_{1},x_{2}\}\times\{y_{1},y_{2}\}$. Let $A,B\subset\Theta$ be two nonempty antichains such that $B>^{\Theta}_{AC}A$. Note that there are three antichains in $\Theta$: $\{(x_{2},y_{2})\},\{(x_{1},y_{1})\},$ and $\{(x_{2},y_{1}),(x_{1},y_{2})\}$. Since only one of these has two elements, at least one of the disjoint antichains $A,B$ is a singleton. Assume w.l.o.g. that this is $B$. Since $B$ hits the maximal value on both coordinates, it must be the case that $B=\{(x_{2},y_{2})\}$. Let $P^{B}$ be the Dirac measure on $B$ and $P^{A}$ be any full support measure on $A$. Then $F^{A}_{1}(x_{1})>0=F^{B}_{1}(x_{1})$ and $F^{A}_{2}(y_{1})>0=F^{B}_{2}(y_{1})$ so $P^{A}\prec^{*}_{cw}P^{B}$.
Inductive Step: Assume the induction hypothesis holds for all composite numbers $s^{\prime}<s$ and consider a state space $\Theta$ with $|\Theta|=s$. Let $A,B\subset\Theta$ be two nonempty antichains such that $B>^{\Theta}_{AC}A$. If one of these is a singleton then the proof is straightforward by an argument similar to the base case. Assume that $|A|,|B|\geq 2$. We can write $A=\{\delta^{1},\ldots,\delta^{k}\},B=\{\theta^{1},\ldots,\theta^{\ell}\}$, where both sets are ordered according to the first coordinate from the least to greatest. That is, $\delta^{1}_{1}<\delta^{2}_{1}<\ldots<\delta^{k}_{1}$, and similarly for $B$. Since $A$ and $B$ are antichains it must be the case that the order on the second coordinate is opposite, i.e., in $A$, $\delta^{1}_{2}>\delta^{2}_{2},\ldots>\delta^{k}_{2}$ and similarly in $B$. Since $A$ hits the minimal value on both coordinates and $B$ hits the maximal value on both coordinates we have $\delta^{1}_{1}\leq\theta^{1}_{1}$ and $\delta^{1}_{2}\leq\theta^{2}_{2},$ so $\delta^{1}\leq\theta^{1}$.
Next, we claim that one of the following holds
$\left\{\begin{array}[]{l}\text{(i) }\quad\delta^{1}\leq\theta^{1},\theta^{2};\text{ or }\\ \text{(ii) }\quad\theta^{1}\geq\delta^{1},\delta^{2}\end{array}\right$. (3)
Assume towards contradiction that this is not the case. Then, since $\delta^{1}$ is minimal on the first coordinate (obtains the minimal value in $\Theta_{1}$), and $\theta^{1}$ is maximal on the second coordinate it must be the case that $\delta^{1}_{2}>\theta^{2}_{2}$ and $\theta^{1}_{1}<\delta^{2}_{1}$. Let $\tau=(\theta^{1}_{1},\max\{\theta^{2}_{2},\delta^{2}_{2}\})\in\Theta$. Note that $\tau\ll\bar{\theta}$.
We claim that for every $\theta\in A\cup B$ neither $\theta\leq\tau$ nor $\theta\gg\tau$ hold, so that $A\cup B$ is compensatory in contradiction to the third condition in Definition 7. To see this, note that $\delta^{1}_{1}\leq\theta^{1}_{1}=\tau_{1}$ and $\delta^{1}_{2}>\max\{\theta^{2}_{2},\delta^{2}_{2}\}=\tau_{2}$. So $\delta^{1}$ is neither weakly below $\tau$ nor strictly above it. Since $A$ is an antichain ordered according to the first coordinate, for any $2\leq r\leq k$ we have $\delta^{r}_{1}\geq\delta^{2}_{1}>\theta^{1}_{1}$ and $\delta^{r}_{2}\leq\delta^{2}_{2}\leq\max\{\theta^{2}_{2},\delta^{2}_{2}\}$, so that all states in $A$ are neither weakly below $\tau$ nor strictly above it.
Likewise, $\theta^{1}_{1}=\tau_{1}$ and, since $\theta^{1}$ is maximal on the second coordinate, $\theta^{1}_{2}>\tau_{2}$. So $\theta^{1}$ is neither weakly below $\tau$ nor strictly above it. Since $B$ is an antichain ordered according to the first coordinate, for any $2\leq r\leq\ell$ we have $\theta^{r}_{1}>\theta^{1}_{1}=\tau_{1}$ and $\theta^{r}_{2}\leq\theta^{2}_{2}\leq\tau_{2}$, so that all states in $B$ are neither weakly below $\tau$ nor strictly above it. Thus, $A\cup B$ is compensatory, a contradiction.
We have established that one of the conditions in (3) holds. Assume w.l.o.g. that (i) holds (the proof for (ii) is symmetric). Let $B^{\prime}=B\setminus\{\theta^{1}\}$ and let $\Theta^{\prime}=\{\theta\in\Theta|\theta_{2}\leq\theta^{2}_{2}\}$. We claim that $B^{\prime}>^{\Theta^{\prime}}_{AC}A$. Since $\delta^{1}_{2}\leq\theta^{2}_{2}$ we have $A\subset\Theta^{\prime}$ and therefore $A$ and $B^{\prime}$ are disjoint antichains in $\Theta^{\prime}$, so the first condition in Definition 7 is satisfied. For the second condition, note that $A$ hits the minimal values in $\Theta^{\prime}$ on both coordinates because these are the same minimal values as in $\Theta$. And $B^{\prime}$ hits the maximal values in $\Theta^{\prime}$ on both coordinates, because the maximal value in $\Theta^{\prime}$ on the first coordinate is the same as in $\Theta$ (and by assumption this is the value of $\theta^{\ell}_{1}$) and the maximal value on the second coordinate is $\theta^{2}_{2}$.
To see that that the third condition in Definition 7 holds, assume towards contradiction that $A\cup B^{\prime}$ is compensatory (in $\Theta^{\prime}$). Let $\bar{\theta}^{\prime}=(\bar{\theta}_{1},\theta^{2}_{2})$ denote the maximal value in $\Theta^{\prime}$. Let $\theta^{*}\in\Theta^{\prime}$ be such that $\theta^{*}\ll\bar{\theta}^{\prime}$ and such that every state in $A\cup B^{\prime}$ is neither weakly below $\theta^{*}$ nor strictly above it. This implies that $\theta^{2}_{1}\leq\theta^{*}_{1}$ (otherwise $\theta^{2}\gg\theta^{*}$), which in turn implies that $\theta^{1}_{1}<\theta^{*}_{1}$. But we also have $\theta^{1}_{2}>\theta^{*}_{2}$, so $\theta^{1}$ is neither weakly below $\theta^{*}$ nor strictly above it. But then every state in $A\cup B$ is neither weakly below $\theta^{*}$ nor strictly above it and $\theta^{*}\ll\bar{\theta}$, so $A\cup B$ is compensatory in $\Theta$ in contradiction to $B>^{\Theta}_{AC}A$.
We have established that $B^{\prime}>^{\Theta^{\prime}}_{AC}A$. Since $|\Theta^{\prime}|<s$, by the induction hypothesis there exist distributions $P^{A},P^{B^{\prime}}$ supported on $A$ and $B^{\prime}$, respectively, such that $P^{A}\prec^{*}_{cw}P^{B^{\prime}}$ on $\Theta^{\prime}$. Recall that $\underaccent{\bar}{\theta}$ is the minimal state in $\Theta$ and $\bar{\theta}^{\prime}=(\bar{\theta}_{1},\theta^{2}_{2})$ is the maximal state in $\Theta^{\prime}$. So $P^{A}\prec^{*}_{cw}P^{B^{\prime}}$ implies that there exists $\varepsilon>0$ such that for every coordinate $i$ and for every $x\in[\underaccent{\bar}{\theta}_{i},\bar{\theta}^{\prime}_{i})$ we have $F^{A}_{i}(x)>F^{B^{\prime}}_{i}(x)+2\varepsilon$. Let $D_{\theta^{1}}$ be the Dirac measure on $\theta^{1}$ and define the mixture distribution $P^{B}=(1-\varepsilon)P^{B^{\prime}}+\varepsilon D_{\theta^{1}}$. Then, for any $x\in[\underaccent{\bar}{\theta}_{1},\bar{\theta}_{1})=[\underaccent{\bar}{\theta}_{1},\bar{\theta}^{\prime}_{1})$,
$F^{B}_{1}(x)\leq(1-\varepsilon)F^{B^{\prime}}_{1}(x)+\varepsilon<F^{B^{\prime}}_{1}(x)+2\varepsilon<F^{A}_{1}(x)$
So, $P^{A}_{1}\prec^{*}_{st}P^{B}_{1}$ in $\Theta_{1}$.
Moving on to the second coordinate, for any $x\in[\underaccent{\bar}{\theta}_{2},\bar{\theta}_{2})$ we have, [^26] The argument for the second coordinate is different because the support of $P^{B}_{2}$ is not equal to the support of $P^{B^{\prime}}_{2}$. So, for values $x\in[\bar{\theta}^{\prime}_{2},\bar{\theta}_{2})$ we don’t have $F^{B^{\prime}}_{2}(x)>F^{A}_{2}(x)$ but rather $F^{B^{\prime}}_{2}(x)=1=F^{A}_{2}(x)$. This weak inequality suffices to get a strict inequality between $F^{B}_{2}$ and $F^{A}_{2}$ since the construction of $P^{B}_{2}$ from $P^{B^{\prime}}_{2}$ involved shifting mass from smaller elements to the maximal element in the new support.
$F^{B}_{2}(x)=(1-\varepsilon)F^{B^{\prime}}_{2}(x)<F^{B^{\prime}}_{2}(x)\leq F^{A}_{2}(x)$
Here the first equality follows from the definition of $P^{B}$ and the fact that $\bar{\theta}_{2}=\theta^{1}_{2}$. So, $P^{A}_{2}\prec^{*}_{st}P^{B}_{2}$ on $\Theta_{2}$ and therefore $P^{A}\prec^{*}_{cw}P^{B}$ on $\Theta$, which completes the induction argument.
Proof of Lemma 7. We prove these in order, by showing that each pair of sets satisfies the three conditions in Definition 7.
Proving that $\min(\Gamma)<^{\Theta}_{AC}\max(\Gamma)$.
  1. 1.
    $\min(\Gamma)$ and $\max(\Gamma)$ are disjoint: Assume towards contradiction that there exists $\theta\in\min(\Gamma)\cap\max(\Gamma)$. There must be one coordinate in which $\theta$ does not obtain the maximal value and one coordinate in which $\theta$ does not obtain the minimal value. Otherwise, if for example $\theta=\bar{\theta}$ then since $\theta\in\min(\Gamma)$ it must be the case that $\Gamma=\{\theta\}$ but then $\Gamma$ is not spanning. Assume w.l.o.g. that $\theta$ is not minimal on the first coordinate and is not maximal on the second coordinate.
    Recall that $\Theta_{1}=\{x_{1},\ldots,x_{m}\}$. We can assume that $x_{1}<x_{2}<\ldots<x_{m}$. The assumption that $\theta$ is not minimal on the first coordinate implies $\theta_{1}=x_{j}$ for some $1<j\leq m$. Define $\theta^{*}=(x_{j-1},\theta_{2})$. Then $\theta^{*}\ll\bar{\theta}$. We will show that for every $\theta^{\prime}\in\Gamma$ it is not the case that $\theta^{\prime}\leq\theta^{*}$ and it is not the case that $\theta^{\prime}\gg\theta^{*}$ so $\Gamma$ is compensatory in contradiction to the third condition in Theorem 5.
    If there exists $\theta^{\prime}\in\Gamma$ with $\theta^{\prime}\leq\theta^{*}$ then $\theta^{\prime}\leq\theta$ and $\theta^{\prime}\neq\theta$ in contradiction to $\theta\in\min(\Gamma)$. If there exists $\theta^{\prime}\in\Gamma$ with $\theta^{\prime}\gg\theta^{*}$ then $\theta^{\prime}\geq\theta$ and $\theta^{\prime}\neq\theta$ in contradiction to $\theta\in\max(\Gamma)$.
  2. 2.
    $\min(\Gamma)$ hits the minimal value on both coordinates and $\max(\Gamma)$ hits the maximal value on both coordinates: Follows immediately from the assumption that $\Gamma$ is spanning.
  3. 3.
    $\min(\Gamma)\cup\max(\Gamma)$ is non-compensatory: Assume towards contradiction that there exists $\theta\in\Theta$, with $\theta\ll\bar{\theta}$, such that for every $\theta^{\prime}\in\min(\Gamma)\cup\max(\Gamma)$ it is not the case that $\theta^{\prime}\leq\theta$ and it is not the case that $\theta^{\prime}\gg\theta$.
    Then there does not exist $\theta^{\prime}\in\Gamma$ with $\theta^{\prime}\gg\theta$ since that would imply $\theta^{\prime}\in\max(\Gamma)$ and $\theta^{\prime}\gg\theta$. Similarly, there does not exist $\theta^{\prime}\in\Gamma$ with $\theta^{\prime}\leq\theta$ as that would imply $\theta^{\prime}\in\min(\Gamma)$ and $\theta^{\prime}\leq\theta$. So $\Gamma$ is compensatory, a contradiction.
Proving that $\min(\Gamma)<^{\Theta}_{AC}\max(\Gamma^{C})$:
  1. 1.
    $\min(\Gamma)$ and $\max(\Gamma^{C})$ are disjoint: By definition of the complement, $\min(\Gamma)\cap\max(\Gamma^{C})\subset\Gamma\cap\Gamma^{C}=\emptyset$.
  2. 2.
    $\min(\Gamma)$ hits the minimal value on both coordinates and $\max(\Gamma^{C})$ hits the maximal value on both coordinates: Follows immediately from the assumption that $\Gamma$ and $\Gamma^{C}$ are spanning.
  3. 3.
    $\min(\Gamma)\cup\max(\Gamma^{C})$ is non-compensatory: Assume towards contradiction that there exists $\theta\in\Theta$, with $\theta\ll\bar{\theta}$, such that for every $\theta^{\prime}\in\min(\Gamma)\cup\max(\Gamma^{C})$ it is not the case that $\theta^{\prime}\leq\theta$ and it is not the case that $\theta^{\prime}\gg\theta$. Then $\Gamma$ includes all states strictly above $\theta$, because if $\theta^{\prime}\gg\theta$ and $\theta^{\prime}\in\Gamma^{C}$ then $\theta^{\prime}\in\max(\Gamma^{C})$ and $\theta^{\prime}\gg\theta$, a contradiction. Similarly, $\Gamma^{C}$ includes all states weakly below $\theta$, as otherwise there exists $\theta^{\prime}\in\min(\Gamma)$ with $\theta^{\prime}\leq\theta$. Thus, $\Gamma$ is biased upward, in contradiction to the second condition in Theorem 5.
Proving that $\min(\Gamma^{C})<^{\Theta}_{AC}\max(\Gamma)$: The arguments for the first two conditions of Definition 7 are exactly as those in the proof that $\min(\Gamma)<^{\Theta}_{AC}\max(\Gamma^{C})$. For the third condition, assume towards contradiction that there exists $\theta\in\Theta$, with $\theta\gg\underaccent{\bar}{\theta}$, such that for every $\theta^{\prime}\in\min(\Gamma^{C})\cup\max(\Gamma)$ it is not the case that $\theta^{\prime}\geq\theta$ and it is not the case that $\theta^{\prime}\ll\theta$. [^27] Here we are using the equivalent definition for a set to be compensatory that is stated in Footnote 18. Then, $\Gamma$ includes all states strictly below $\theta$, because if there existed $\theta^{\prime}\in\Gamma^{C}$ with $\theta^{\prime}\ll\theta$ then $\theta^{\prime}\in\min(\Gamma^{C})$, a contradiction. Similarly, $\Gamma^{C}$ includes all states weakly above $\theta$. Thus, $\Gamma$ is biased downward in contradiction to the second condition in Theorem 5.
Proof of Lemma 8. Note that $P^{L}|_{\Gamma}=(1-\delta)P^{L}|_{\min(\Gamma)}+\delta U_{\Gamma\setminus\min(\Gamma)}$ and $P^{L}|_{\Gamma^{C}}=(1-\delta)P^{L}|_{\max(\Gamma^{C})}+\delta U_{\Gamma^{C}\setminus\max(\Gamma^{C})}$. Thus, for every $i\in[m-1]$ and for $\delta$ sufficiently small such that $(1-\delta)\varepsilon>\delta$ we have
$\begin{split}F^{L}_{1}|_{\Gamma}(x_{i})&=(1-\delta)F^{L}_{1}|_{\min(\Gamma)}(x_{i})+\delta F^{U}_{1}|_{\Gamma\setminus\min(\Gamma)}(x_{i})\geq(1-\delta)F^{L}_{1}|_{\min(\Gamma)}(x_{i})\\ &>(1-\delta)F^{L}_{1}|_{\max(\Gamma^{C})}(x_{i})+(1-\delta)\varepsilon\geq(1-\delta)F^{L}_{1}|_{\max(\Gamma^{C})}(x_{i})+\delta\cdot 1\\ &\geq(1-\delta)F^{L}_{1}|_{\max(\Gamma^{C})}(x_{i})+\delta F^{U}_{1}|_{\Gamma^{C}\setminus\max(\Gamma^{C})}(x_{i})=F^{L}_{1}|_{\Gamma^{C}}(x_{i})\end{split}$
Similarly, $F^{L}_{2}|_{\Gamma}(y_{j})>F^{L}_{2}|_{\Gamma^{C}}(y_{j})$ for every $j\in[n-1]$. So $P^{L}|_{\Gamma}\prec^{*}_{cw}P^{L}|_{\Gamma^{C}}$. An analogous argument shows that $P^{H}|_{\Gamma}\succ^{*}_{cw}P^{H}|_{\Gamma^{C}}$. Thus, by Lemma 5, for sufficiently small $\delta$ we have $Q^{L}\prec^{*}_{cw}P^{L}$ and $P^{H}\prec^{*}_{cw}Q^{H}$. [^28] This is stronger than we need since strong coordinatewise polarization only requires a strong ordering between the priors.
It remains to prove that $P^{L}\prec^{*}_{cw}P^{H}$. For every $i\in[m-1]$ and for sufficiently small $\delta$,
$\begin{split}F^{L}_{1}(x_{i})&\geq(1-\delta)^{2}F^{L}_{1}|_{\min(\Gamma)}(x_{i})\\ &>(1-\delta)^{2}F^{H}_{1}|_{\max(\Gamma)}(x_{i})+(1-\delta)^{2}\varepsilon>(1-\delta)^{2}F^{H}_{1}|_{\max(\Gamma)}(x_{i})+(1-\delta)2\delta+\delta^{2}\\ &\geq(1-\delta)^{2}F^{H}_{1}|_{\max(\Gamma)}(x_{i})+(1-\delta)\delta F^{U}_{1}|_{\Gamma\setminus\max(\Gamma)}(x_{i})\\ &+(1-\delta)\delta F^{H}_{1}|_{\min(\Gamma^{C})}(x_{i})+\delta^{2}F^{U}_{1}|_{\Gamma^{C}\setminus\min(\Gamma^{C})}(x_{i})=F^{H}_{1}(x_{i})\end{split}$
Here we take $\delta$ such that $(1-\delta)^{2}\varepsilon>(1-\delta)2\delta+\delta^{2}$. This inequality is satisfied for sufficiently small $\delta>0$ because as $\delta\rightarrow 0$ the right-hand-side converges to 0 while the left-hand-side converges to $\varepsilon>0$. A similar argument shows that $F^{L}_{2}(y_{j})>F^{H}_{2}(y_{j})$ for every $j\in[n-1]$. So $P^{L}\prec^{*}_{cw}P^{H}$.